### v1

# Polynomial-like dynamics of analytic maps

Abstract.
The theory of polynomial-like maps is of fundamental importance in holomorphic dynamics. We study dynamical properties of a larger class of maps. Our main result is that, under some natural conditions, a map of this class has a completely invariant compact set if and only if this set is the filled Julia set of a polynomial-like restriction of the map. We also generalize this result to include maps with non-connected domains of definition.
1. Introduction and main results
1.1. Polynomial-like mappings
The theory of polynomial-like mappings was developed by Douady and Hubbard [2] to explain, in particular, the fractal nature of non-linear phenomena in holomorphic dynamics. In this paper, proper holomorphic maps which are not polynomial-like are considered and we are interested in conditions that guarantee a polynomial-like restriction of such a map of the same degree.
Recall that a polynomial-like (PL) map is a triple $(V^{\prime},V,g)$ where $V^{\prime},V$ are topological disks in the plane such that $\overline{V^{\prime}}\subset V$ and $g:V^{\prime}\to V$ is a proper holomorphic map of some degree $d\geq 2$ [2]. The non-escaping set $K_{g}:=\{z:g^{n}(z)\in U_{-1},n=0,1,2,$...\} is called the filled Julia set of the PL map $g$.
It is easy to see that $K_{g}$ is a compact set, it is full, i.e., its complement $\mathbb{C}\setminus K_{g}$ is connected, and is completely invariant, i.e., $g^{-1}(K_{g})=K_{g}$.
By the Straightening Theorem [2], a PL map $g:V^{\prime}\to V$ is hybrid equivalent (i.e., conjugate in some neighborhood of $K_{g}$ by a quasiconformal homeomorphism which is conformal a.e. on $K_{g}$) to a polynomial of degree $d$. This implies that the set of repelling periodic points of the map $g:K_{g}\to K_{g}$ is dense in $\partial K_{g}$, moreover, for every $n\in\mathbb{N}$, the map $g^{n}:K_{g}\to K_{g}$ has $d^{n}$ fixed points (counting multiplicity), every component of the interior of $K_{g}$ is (pre)periodic, and the set of limit points of the set $\{g^{-n}(x)\}_{n\geq 0}$ coincide with $\partial K_{g}$, for all $x\in K_{g}$ with the only exception if $g^{-1}(x)=\{x\}$ (in the latter case, $x$ is the only critical point of $g$, it is fixed by $g$ and has the maximal multiplicity $d$).
In the present paper we study similar maps but give up the condition that the domain of definition is compactly contained in the range. Our main result is that if such a map has a completely invariant set $K$ which is a full compact containing all critical points then the map admits a polynomial-like restriction to a neighborhood of $K$ and $K$ is its filled Julia set, see Theorem 1.1 and Corollary 1.1 for complete claims. In Theorem 1.3 we generalize this result to include maps with non-connected domains of definition.
1.2. Assumptions and notations
Throughout the paper - with the only exception Section 1.4 (see Theorem 1.3 and Remark 1.3) and Section 6 (where Theorem 1.3 is proved) - we fix a proper holomorphic map $g:U^{\prime}\to U$ of some degree $d\geq 2$ as follows:
  • $U^{\prime},U$ are simply connected (perhaps, unbounded) domains in $\mathbb{C}$ such that $U^{\prime}\subset U$ and $U^{\prime}\neq U$,
  • let $C_{g}=\{x:g^{\prime}(x)=0\}$ be the set of all critical points of $g$, then all iterates $g^{n}(c)$, $n\geq 0$, are well-defined for any $c\in C_{g}$ and the postcritical set
    $P_{g}=\overline{\{g^{n}(C_{g}):n=0,1,...\}}$
    is a compact subset of $U^{\prime}$,
  • if $P_{g}$ is a singleton $\{c\}$ then $g^{-1}(c)\neq\{c\}$.
Let $U_{-i}=g^{-i}(U)$, $n=0,1,2,$.... In particular, $U_{0}=U$ and $U_{-1}=U^{\prime}$. Then $\{U_{-i}\}_{i=0}^{\infty}$ is a decreasing sequence of (connected) simply connected domains and $P_{g}\subset\cap_{i\geq 0}U_{-i}$.
Other notations.
  • $BO(x)=\{y:g^{n}(y)=x\mbox{ for some }n\geq 0\}$ is the backward orbit of $x\in U$.
  • $O(x)=\{g^{n}(x):n=0,1,2,$...\}, the forward orbit of $x$ assuming all iterates $g^{i}(x)$, $i\geq 0$ are defined, i.e, $x\in\cap_{i\geq 0}U_{-i}$. In particular, $O(x)$ is defined for any point of a (forward) invariant set $E\subset U^{\prime}$, i.e., such that $g(E)\subset E$). In this case, $\omega(x)$ denotes, as usual, the set of limit points of $O(x)$.
  • Grand orbit of $x$ is
    $GO(x)=\{z:g^{m}(z)=g^{n}(z)\mbox{ for some }m,n\geq 0\},$
    assuming that $O(x)$ is well-defined; in particular, $GO(c)$ is defined for each $c\in C_{g}$,
  • if $X\subset\mathbb{C}$ is compact, let $\hat{X}$ (the topological hull of $X$) be the complement to the unbounded component of $\mathbb{C}\setminus X$. Then $\hat{X}$ is a full compact. Note that if $X\subset\Omega$ where $\Omega$ is simply connected, then $\hat{X}\subset\Omega$.
1.3. Main results
Theorem 1.1.
[…]
Corollary 1.1.
For the map $g:U^{\prime}\to U$, there exists a unique maximal (by inclusion) full completely invariant compact $K_{*}\subset U$ and one and only one of the following alternatives holds:
  1. (1)
    $K_{*}=\emptyset$,
  2. (2)
    $P_{g}\subset K_{*}$ and $K_{*}$ is the filled Julia set of a PL restriction as in Theorem 1.1; moreover, given $x\in U\setminus K_{*}$, either $g^{n}(x)\in U\setminus U^{\prime}$ for some $n\geq 0$, or the forward orbit $O(x)$ is well-defined but
    $\omega(x)\subset\partial U,$
  3. (3)
    $\partial K_{*}\subset P_{g}$ and $C_{g}\setminus K_{*}\neq\emptyset$; in fact, $K_{*}=\{z:GO(z)\subset\hat{P}_{g}\}$,
Furthermore, if $Z$ is any completely invariant compact set (not necessarily full) of $g:U^{\prime}\to U$, then $Z\subset K_{*}$.
Additionally, let $R_{g}$ be the closure (in the Riemann sphere $\hat{\mathbb{C}}$) of the set of repelling periodic points of $g:U^{\prime}\to U$. Then $R_{g}=\partial K_{*}$ in the case (2) while in cases (1) and (3) the compact $R_{g}$ (in $\hat{\mathbb{C}}$) contains a (non trivial) continuum which has common points with the range $U$ as well as with its boundary $\partial U$.
Note that $K_{*}\subset\cap_{i\geq 0}U_{-i}$ and $x\in K_{*}$ if and only if $GO(x)\subset K_{*}$. In particular, $C_{g}\setminus K_{*}\neq\emptyset$ holds if and only if $P_{g}\nsubseteq K_{*}$.
See Sect 5 for the proof of the corollary.
Remark 1.2.
It is not clear whether the alternative (3) can occur. Our proof of Theorem 1.1 uses the fact that $P_{g}\subset K$.
The proof of Theorem 1.1 is a mixture of dynamical and geometric arguments. First, we show that either $g:U^{\prime}\to U$ admits a PL restriction as in Theorem 1.1 or it has a lot of repelling periodic orbits close to $\partial U$, see Proposition 2.2 and Corollary 2.3. Therefore, Theorem 1.1 will be an immediately consequence of the next
Theorem 1.2.
Under the conditions of Theorem 1.1, the map $g:U^{\prime}\to U$ has no repelling periodic points outside of $K$.
The proof of Theorem 1.2 in the case of connected $K$ is rather straightforward: using the Riemann map $\pi:\{|z|>1\}\to\mathbb{C}\setminus K$ one can conjugate the map $g:U^{\prime}\setminus K\to U\setminus K$ to an unbranched $d$-covering map in a one-sided neighborhood of the unit circle $S^{1}$ and apply Schwartz’s reflection principle to extend it through $S^{1}$; then one can end the proof, for example, by noticing that a repelling cycle outside $S^{1}$ would attract inverse iterates of points of $S^{1}$, in contradiction with the complete invariance of $S^{1}$. It is not clear how to generalize this idea to the disconnected $K$. We propose another proof in the case of connected $K$, which is based on a different idea, see Sect. 3.1, and then carry this idea over to the disconnected case. The latter part is more technical and uses general results of the ergodic theory of locally analytic maps and the uniformization theory of plane sets, see Appendices A-B.
1.4. A generalization
Note that a priori the completely invariant compact $K$ of Theorem 1.1 is not required to be connected although a posteriori $K$ is connected because it contains (iterates of) all critical points of the PL map $g:V^{\prime}\to V$.
We have the following generalization of Theorem 1.1 where $K$ is always disconnected.
Remind that a generalized PL map [4] $G:V^{\prime}\to V$ is a holomorphic map between open sets $V^{\prime},V$, where $V$ is a topological disk, $\overline{V^{\prime}}\subset V$, and $V^{\prime}$ consists of a finitely many simply connected components $V_{k}^{\prime}$, k=1,,Mk=1,...,M, with pairwise disjoint closures, such that $G:V^{\prime}\to V$ is a proper map of some degree $D\geq 2$. Its set of non-escaping points $K_{G}=\{z:G^{n}(z)\in V^{\prime},n=0,1,$...\} - the filled Julia set of $G$ - is a completely invariant full compact. (Note that if $V^{\prime}$ is connected, $G:V^{\prime}\to V$ is the classical Douady-Hubbard PL map.) The Straightening theorem also holds: $G:V^{\prime}\to V$ is conjugated, in a neighborhood of $K_{G}$, to a polynomial of degree $D$ (see e.g. [3] for a proof of a similar claim). If $M\geq 2$, then $K_{G}$ is never connected.
Theorem 1.3.
Suppose that $U^{\prime}\subset U$ are open subsets of $\mathbb{C}$ where $U$ is connected simply connected and $U^{\prime}$ consists of a finitely many simply connected components $U^{\prime}_{k}$, k=1,,mk=1,...,m, for some $m\geq 2$. Let $g:U^{\prime}\to U$ be a proper holomorphic map of some degree $d\geq 2$. Let $K\subset U$ be a full compact set such that $g^{-1}(K)=K$ and $K$ contains all critical points of $g:U^{\prime}\to U$ (if any).
Then there exist a topological disk $V_{0}$, $\overline{V_{0}}\subset U$, and $N\geq 1$ such that $\overline{V_{-N}}\subset V_{0}$, where $V_{-N}=g^{-N}(V_{0})$, and $g^{N}:V_{-N}\to V_{0}$ is a generalized PL map of degree $d^{N}$ and $K$ is its filled Julia set.
Moreover, there exists a neighborhood $V\subset V_{0}$ of $K$ consisting of a finitely many simply connected components as follows: if $V^{\prime}:=g^{-1}(V)$, then $\overline{V^{\prime}}\subset V$ and $g:V^{\prime}\to V$ is a proper map of degree $d$, such that $K=\{x:g^{n}(x)\in V^{\prime},n=0,1,2,$...\}.
Remark 1.3.
Let again $g:U^{\prime}\to U$ be a proper holomorphic map of some degree $d\geq 2$ where $U^{\prime},U$ are simply connected domains, and $K$ is a full compact which is completely invariant. Suppose that not all critical points of $g$ are in $K$ and that any critical point $c$ of $g$ outside of $K$ escapes, i.e., $g^{n}(c)\in U\setminus U^{\prime}$ for some $n=n(c)>0$. Then Theorem 1.3 applies to some iterate $g^{M}:g^{-M}(U)\to U$ of $g$ (with e.g. $M=\max\{n(c)\}$) and the compact $K$.
Acknowledgments. I would like to thank Feliks Przytycki for several fruitful discussions concerning mainly Theorem A.1, and Misha Sodin, Sebastian van Strien and Vladlen Timorin for helpful comments and references.
2. Preliminary statements
Lemma 2.1.
Basin of every attracting or parabolic cycle of $g:U^{\prime}\to U$ contains a critical point of $g$.
Proof.
The proof repeats the classical one: assume that the immediate basin $B$ of some attracting or parabolic cycle $P$ of $g$ contains no critical values of $g$. Then the inverse of Konigs (if $P$ is attracting and not superattracting) or Fatou (if $P$ is parabolic) linearizing coordinate extends, by the inverse branches of $g$, to an entire function with values in $B$, a contradiction with Liouville’s theorem. ∎
Next, we are going to describe $3$ alternatives. Let us fix a Jordan curve $\Gamma_{0}$ in $U$ which encloses the postcritical set $P_{g}$. Let $V_{0}$ be the Jordan domain bounded by $\Gamma_{0}$, so that $P_{g}\subset V_{0}\subset\overline{V_{0}}\subset U$.
Let, for each $n>0$, $V_{-n}=g^{-n}(V_{0})$ (which is compactly contained in $g^{-n}(U)$) and $\Gamma_{n}=\partial V_{-n}$ As $P_{g}\subset V_{0}$, then $V_{-n}$ is also simply connected and $\Gamma_{n}$ is a Jordan curve.
A priori, there are $3$ possibilities:
(A) $\Gamma_{n}\cap\Gamma_{0}\neq\emptyset$ for all $n$.
(B) For some $n>0$, $\overline{V}_{-n}\subset V_{0}$.
(C) For some $n>0$, $\overline{V}_{0}\subset V_{-n}$.
Proposition 2.2.
Assume that (A) holds. Then, for every $\varepsilon>0$ and every $n$ big enough, the map $g^{n}$ has a repelling fixed point in $\varepsilon$-neighborhood of $\Gamma_{0}$.
Assume that (B) holds. Then $g^{n}:V_{-n}\to V_{0}$ is a PL map of degree $d^{n}$. Furthermore, if there exists a full compact $K\subset V_{0}$ such that $g^{-1}(K)=K$, then $K=K_{G}=\cap_{k\geq 0}g^{-kn}(V)$, the filled Julia set of the PL map $g^{n}:V_{-n}\to V_{0}$. Moreover, there exists a polynomial-like restriction $g:V^{\prime}\to V$ where $V$ is a simply connected neighborhood of $K$ and $K$ is the filled Julia set of $g:V^{\prime}\to V$. In particular, $P_{g}\subset K$
Assume that (C) holds. Then the restricted map $g:\tilde{V}^{\prime}\to\tilde{V}$ where $\tilde{V}^{\prime}=\cup_{i\geq 1}g^{-i}(V_{0})$, $\tilde{V}=\cup_{i\geq 0}g^{-i}(V_{0})$ is trivial: there is an attracting fixed point $a\in\cap_{i\geq 0}g^{-i}(V_{0})$ such that $g^{n}(z)\to a$ uniformly in $V$.
Moreover, if $\Gamma_{0}$ is close enough to $\partial U$, then (C) does not happen.
Proof.
(A). In this case, there exist a sequence of indices $n_{i}\to\infty$ and a sequences of points $x_{i}\in\Gamma_{n_{i}}\cap\Gamma_{0}$ as follows: $x_{i}\to b^{\prime}$, $y_{i}:=g^{n_{i}}(x_{i})\in\Gamma_{0}$, $y_{i}\to b$ and $(g^{n_{i}})^{\prime}(x_{i})\to L$ where $L\in\mathbb{C}\cup\{\infty\}$. Fix a small disk $B$ centered at $b$ which is contained in $U$ and disjoint with $P_{g}$, and, for each $i$, let $\phi_{i}$ be a branch of $g^{-n_{i}}$ which is well-defined in $B$ and maps $y_{i}$ to $x_{i}$. Then $(\phi_{i})^{\prime}(y_{i})\to 1/L$.
Thus $\phi_{i}:B\to\mathbb{C}$, $i=1,2,$..., is a sequence of univalent maps and is a normal family because it avoids taking values in the set $P_{g}\cup g^{-1}(P_{g})$, which must contain at least $2$ points. By this reason, the case $L=0$ is impossible.
Consider the case $L\neq\infty$, that is, $0<|L|<\infty$. Replacing $(n_{i})$ by a subsequence, in this case $(\phi_{i})$ tends uniformly on compacts in $B$ to a non-constant, hence, univalent function $\nu$ which maps $B$ to a neighborhood $B^{\prime}$ of $b^{\prime}$. As $\psi_{i}(B)\subset g^{-n_{i}}(U)$, then $B^{\prime}\subset\cap_{i\geq 0}g^{-i}(U)$. Shrinking slightly $B^{\prime}$, we get:
Claim. The sequence $g^{n_{i}}:B^{\prime}\to B$ is well-defined and converges uniformly to $\nu^{-1}$.
Let $m_{i}=n_{n+1}-n_{n_{i}}$ Passing again to a subsequence of $(n_{i})$, one can assume that $m_{i}\to\infty$. Write $g^{n_{i+1}}=g^{m_{i}}\circ g^{n_{i}}$, on $B^{\prime}$. Here, $g^{n_{i}}(B^{\prime})\subset g^{n_{i}}(g^{-n_{i+1}}(U))\subset g^{-m_{i}}(U)$.
Shrinking again $B^{\prime}$ if necessary, by the Claim, $g^{m_{i}}\to\operatorname{Id}$, uniformly on $\nu^{-1}(B^{\prime})$ where $\nu^{-1}(B^{\prime})\subset g^{-m_{i}}(U)$. Concluding this argument (and shrinking $B$ to a smaller disk around $b$), we obtain that, for some branches $I_{i}:=g^{-m_{i}}$ on $B$, $I_{i}\to\operatorname{Id}$ uniformly on $B\subset U_{0}$.
Let now $D$ be any simply connected subdomain of $U\setminus P_{g}$ containing $B$. As all branches of $g^{-m_{i}}$ are well-defined in $D$ and form a normal family, passing to a subsequence, we get that $I_{i}$ (more precisely, their extensions from $B$ to $D$) converge uniformly on compacts in $D$ to $\operatorname{Id}$. As each $I_{i}$ is an inverse to $g^{m_{i}}$, it follows that, given a compactly contained in $D$ domain $F\subset D$, for all $i$ big enough, we have: $F\subset g^{-m_{i}}(U)$ and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $F$.
We apply this to two domains $D_{1},D_{2}$ which together cover the curve $\Gamma_{0}$, and to their compactly contained subdomains $F_{1}\subset D_{1}$, $F_{2}\subset D_{2}$, such that $F_{1}\cup F_{2}\supset\Gamma_{0}$. As a result, we obtain that $\Gamma_{0}\subset\cap_{i\geq 0}g^{-m_{i}}(U)$ for all $i$ big enough and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $\Gamma_{0}$. This is a contradiction by the following reason: on the one hand, as we have just shown, $\Gamma_{0}\subset g^{-m_{i}}(U)$, also $\Gamma_{0}=\partial V_{0}$ and $g^{m_{i}}(g^{-m_{i}}(U))=U$, hence, the sequence $g^{m_{i}}$ is well-defined on $\overline{V_{0}}$ and converges uniformly to the identity map $\operatorname{Id}$. Then $(g^{m_{i}})^{\prime}\to 1$ uniformly on compacts in $V_{0}$. On the other hand, $(g^{m_{i}})^{\prime}(c)=0$ for each $c\in C_{g}$ This contradiction excludes the case $0<|L|<\infty$.
Thus we must have that $(g^{n_{i}})^{\prime}(x_{i})\to\infty$, i.e., $(\psi_{i})^{\prime}(y_{i})\to 0$. Let us show that then $g^{n_{i}}$ must have a repelling fixed point close to the curve $\Gamma_{0}$. Indeed, let $D\subset U_{0}\setminus P_{g}$ be a simply connected domain that contains points $b^{\prime}=\lim x_{i}$ and $b=\lim y_{i}$. Then, by the normality of the family of extensions of $\phi_{i}$ from $B$ to $D$, a subsequence of $\phi_{i}$ converges uniformly on compacts in $D$ to a point $A\in D$. In particular, for a small disk $B_{A}$ centered at $A$, and for some branch $\phi_{i}$ of $g^{-n_{i}}$, $\overline{\phi_{i}(B_{A})}\subset B_{A}$. Therefore, $\phi_{i}$ has an attracting fixed point $a\in B_{A}$, which is thus a repelling fixed point of $g^{n_{i}}$.
(B). $\overline{V}_{-n}\subset V_{0}$ for some fixed $n>0$. Then $G:=g^{n}:V_{-n}\to V_{0}$ is a polynomial-like map (of degree $d^{n}$). Let $K_{G}$ be the filled Julia set of $G$. Then $K_{G}=\cap_{i\geq 0}G^{-i}(V_{0})$.
Assume now that additionally there exists a compact $K$ as in the condition, i.e., $K\subset V_{0}$ and $K$ is full and completely invariant. This implies that $G^{-1}(K)=K$ and that $K\subset\cap_{i\geq 0}G^{-i}(V_{0})=K_{G}$. As $G$ is conjugate to a polynomial near $K_{G}$, it follows that $K=K_{G}$. Moreover, as $P_{G}=P_{g}\subset V_{0}$, hence, $K=K_{G}$ is connected.
It remains to show $g$ itself has a PL restriction, i.e., there exist a simply connected neighborhood $V$ of $K$, such that $V^{\prime}:=g^{-1}(V)$ is compactly contained in $V$. If $n=1$, there is nothing to prove (set $V=V_{0}$). So let $n\geq 2$.
We present two proofs: the first one uses heavily the fact that $K$ is a continuum while the second one holds for non-connected $K$ as well and can be easily adapted to the set up of Theorem 1.3 (where $K$ is disconnected).
First proof.
The following argument is well-known. Let $\Psi:\mathbb{C}\setminus K\to\mathbb{D}^{*}:=\{|w|>1\}$ be the conformal isomorphism such that $\Psi(\infty)=\infty$. Below, $\hat{A}=\Psi(A\setminus K)$ for a neighborhood $A$ of $K$, and $A_{*}$ is the union of $\hat{A}$ with its symmetric w.r.t. $S^{1}$ completed by $S^{1}$. Similar notations are used for maps. Replace $g:U_{-1}\setminus K\to U_{0}\setminus K$ by $\hat{g}:=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-1}\to\hat{U}_{0}$. By Schwartz’s reflexion principle, $\hat{g}$ extends to a holomorphic map $g_{*}:(U_{-1})_{*}\to(U_{0})_{*}$. The latter map has no critical points as $g_{*}^{-1}(S^{1})=S^{1}$. Let $\hat{G}=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-m}\to\hat{U}_{0}$ and $G_{*}:(U_{-m})_{*}\to(U_{0})_{*}$. As this map is a covering and $\overline{(U_{-m})_{*}}\subset(U_{0})_{*}$, then $G_{*}$ expands the hyperbolic metric $\rho$ of the domain $(U_{0})_{*}$, i.e, $D_{G_{*}}(w):=\rho\circ G_{*}(w)|G_{*}^{\prime}(w)|/\rho(w)>1$ for $w\in(U_{-m})_{*}$. Now note that $G_{*}=g_{*}^{m}$ in some neighborhood of $S^{1}$. Define a new metric $\tilde{\rho}=\sum_{i=0}^{m-1}\rho\circ g_{*}^{i}|(g_{*}^{i})^{\prime}|$. Then $\tilde{D}_{g_{*}}(w):=\frac{\tilde{\rho}\circ g_{*}(w)|g_{*}^{\prime}(w)|}{\tilde{\rho}(w)}=1+\frac{\rho(w)}{\tilde{\rho}(w)}\{D_{G_{*}}(w)-1\}>1$ for all $w$ near $S^{1}$. By the compactness, $\tilde{D}_{g_{*}}(w)\geq\lambda$ for some $\lambda>1$ and all $w$ in a small annulus around $S^{1}$. It follows that there is a neighborhood $V_{*}$ of $S^{1}$ which is symmetric w.r.t. $S^{1}$ such that $g_{*}^{-1}(V_{*})$ is compactly contained in $V_{*}$. If we let $V=\Psi^{-1}(V_{*}\cap\mathbb{D}^{*})\cup K$ then $V$ is a simply connected neighborhood of $K$ such that $\overline{V^{\prime}}\subset V$ where $V^{\prime}=g^{-1}(V)$.
Second proof. Here we don’t use that $K$ is a continuum. For every $j\geq 0$, let $\Omega_{j}:=\cup_{i=j}^{j+n-1}g^{-i}(V_{0})$. As $V_{-n}=g^{-n}(V_{0})$ and $\overline{V_{-n}}\subset V_{0}$, then $K=K_{G}=\cap_{i>0}g^{-ni}(V_{0})$ Hence, by continuity, there exists $j_{*}>0$ such that, for every $j\geq j_{*}$, $K\subset\Omega_{j}\subset V_{0}$ and $\Omega_{j}$ shrink to $K$ as $j\to\infty$. This allows us to write, for every $j\geq j_{*}$:
$g^{-1}(\Omega_{j})=\cup_{i=j+1}^{j+n}g^{-i}(V_{0})=\cup_{i=j+1}^{j+n-1}g^{-i}(V_{0})\cup g^{-j-n}(V_{0})=\Omega_{j}\setminus g^{-j}(V_{0})\cup g^{-j-n}(V_{0}).$
Therefore, $g^{-1}(\Omega_{j})\subset\Omega_{j}$ and
$g:g^{-1}(\Omega_{j})\setminus K\to\Omega_{j}\setminus K$
is an unbranched covering map, for each $j\geq j_{*}$.
Let $\Omega^{*}=\Omega_{j_{*}}\setminus K$ and let $\rho_{*}$ be the hyperbolic metric of the domain $\Omega^{*}$. As $g^{-1}(\Omega^{*})\subset\Omega^{*}$ and there is no an equality, by the above and the Schwartz lemma, $g^{-1}:\Omega^{*}\to\Omega^{*}$ is a local strict contraction.
Now, fix $j$ large enough, such that $\overline{\Omega_{j}}\subset\Omega_{j_{*}}$.
Let $Y$ be the topological hull of $\overline{\Omega_{j}}$. As $g^{-1}(\overline{\Omega_{j}})\subset\overline{\Omega_{j}}$ and $g:U^{\prime}\to U$ is a proper map between simply connected domains, $g^{-1}(Y)\subset Y$. Hence, if $Y^{*}=Y\setminus K$, then $g^{-1}(Y^{*})\subset Y^{*}$ as well (remind that $K$ is completely invariant).
The set $\Delta:=Y\setminus Int(g^{-1}(Y))$ is a compact subset of $\Omega^{*}$, therefore, there exist $r>0$ and $\kappa\in(0,1)$ as follows: for every $a\in\Delta$, $B(a,2r)$ is a topological disk and every branch $g^{-1}:B(a,2r)\to\Omega_{*}$ contracts the distance in $\Omega^{*}$ by a factor at leat $\kappa$ (here and below in the proof $B(x,t)$ is the disk centered at $x\in\Omega^{*}$ and of radius $t>0$ in the hyperbolic metric $\rho_{*}$ of $\Omega^{*}$).
Note that $\partial Y\subset\partial\Omega_{j}\subset\Omega^{*}$ and $\partial Y^{*}$ is the disjoint union of $\partial Y$ and $\partial K$. Let $Y^{*}_{r}=Y^{*}\cup_{a\in\partial Y}\bar{B}(a,r)$.
Let $\Gamma_{*}$ be the outer boundary of $Y^{*}_{r}$ (i.e, the boundary of the topological hull of $Y^{*}_{r}$). Note that $\Gamma_{*}$ is a compact subset of $\Omega^{*}$. Let $V_{*}$ be a (simply connected) domain bounded by $\Gamma_{*}$. We claim that $g^{-1}(V_{*})$ is compactly contained in $V_{*}$. It’s enough to show that $g^{-1}(\Gamma_{*})\subset V_{*}$. To this end, let $z\in\Gamma_{*}$ and $z^{\prime}\in g^{-1}(z)$. There exist $a\in\partial Y$, such that $\rho_{*}(z,a)\leq r$, and a branch $\psi$ of $g^{-1}$ in a neighborhood of $\bar{B}(a,r)$ such that $\psi(z)=z^{\prime}$. If $a^{\prime}:=\psi(a)$, then $\rho_{*}(a^{\prime},z^{\prime})\leq k\rho_{*}(a,z)\leq kr<r$. As $a^{\prime}\in g^{-1}(\partial Y)\subset\Delta$, it follows that $z^{\prime}\in Y^{*}_{r}$. Thus $V^{\prime}_{*}:=g^{-1}(V_{*})$ is compactly in $V_{*}$, i.e., $g:V_{*}^{\prime}\to V*$ is the required PL map.
(C). If $\overline{V}_{0}\subset V_{-n}$ for some $n>0$, then $V_{0}\subset g^{-n}(V_{0})\subset g^{-n}(U)\subset U^{\prime}$. Therefore, if (C) holds for every $V_{0}$, then $U\subset U^{\prime}\subset U$, i.e., $U=U^{\prime}$, a contradiction. This proves the last claim, i.e., if $\Gamma_{0}$ is close enough to $\partial U$, then (C) don’t happen. For the first claim, notice that $\overline{g^{n}(V_{0})}\subset V_{0}$. Hence, by the Wolff-Denjoy theorem (eg [1], the sequence of domains $g^{kn}(V_{0})$, $k=1,2,$..., shrink to an attracting fixed point $a\in V_{0}$ of $g^{n}$. Then $g^{i}(\overline{V_{0}})$ tend to the cycle $O\ni a$ such that $O\subset P_{g}\subset V_{0}$. Therefore, $O=g^{i}(O)\subset g^{i}(V_{0})$ for all $i\geq 0$. As $g^{i}(V_{0})\to O$, it follows that $O=\{a\}$.
We have the following
Corollary 2.3.
Assume that $g:U^{\prime}\to U$ has no a PL restriction to a neighborhood of its postcritical set $P_{g}$. Let $R_{g}$ be the closure (in the Riemann sphere $\hat{\mathbb{C}}$) of the set of repelling periodic points of $g:U^{\prime}\to U$. Then $R_{g}=\partial K_{*}$ in the case (2) while in cases (1) and (3) the compact $R_{g}\subset\hat{\mathbb{C}}$ contains a (non trivial) continuum which has common points with the range $U$ as well as with its boundary $\partial U$.
Proof.
[…]
Repeating an argument as in the case (A) of the proof above, we obtain
Lemma 2.4.
Let $a$ be a repelling periodic point of $g$ in the unbounded component $\Omega$ of $\mathbb{C}\setminus P_{g}$. If $z\in U\cap\Omega$, then for every simply connected domain $W$ in $\Omega$ that contains points $a$ and $z$ there exists a sequence of inverse branches $g^{-n_{k}}:W\to\mathbb{C}$ such that $g^{-n_{k}}(W)\to a$ uniformly on compacts.
Next claim describes properties of completely invariant sets inside of a forward invariant one.
Proposition 2.5.
Let $X$ be a compact forward invariant set such that $P_{g}\subset X$. Let
$D_{\infty}=U\setminus\hat{X}$,
where $\hat{X}$ is the topological hull of $X$. Then:
  1. (1)
    $g^{-1}(D_{\infty})\subset D_{\infty}$,
  2. (2)
    $g(\hat{X})\subset\hat{X}$.
If
$X_{*}=\{x\in\hat{X}:GO(x)\subset\hat{X}\}$
then
  1. (3)
    $X_{*}$ is the maximal completely invariant compact (perhaps, empty) which is contained in $\hat{X}$. Furthermore, $X_{*}$ is a full compact.
Finally, assuming $X_{*}\neq\emptyset$, either $P_{g}\subset X_{*}$ or $\partial X_{*}\subset P_{g}$.
Proof.
We split it into several parts.
Step I. We show that:
(1) $g^{-1}(D_{\infty})\subset D_{\infty}$.
(2) $g(\hat{X})\subset\hat{X}$.
Proof of (1). Since $X$ is forward invariant, then $X\subset\cap_{i\geq 0}U_{-i}$. It is enough to prove that, for every $i\geq 0$, $g_{i}^{-1}(U_{-i}\setminus\hat{X})\subset U_{-i-1}\setminus\hat{X}$. As $U_{-i}\setminus\hat{X}$ is a connected open set, one can join any $z\in U_{-i}\setminus\hat{X}$ with $\partial U_{-i}$ by a simple curve $\gamma$ in $U_{-i}\setminus\hat{X}$. Let $\gamma^{\prime}$ be a component of $g_{i}^{-1}(\gamma)$. It is also a simple curve that joins a preimage of $z$ to $\partial U_{-i-1}$. If $\gamma^{\prime}$ crossed $\partial\hat{X}$ then $\gamma$ would contain a point of $\hat{X}$ because $g(\partial\hat{X})\subset g(X)\subset X\subset\hat{X}$, a contradiction. Therefore, $\gamma^{\prime}\subset U_{-i-1}\setminus\hat{X}$.
(2) follows from (1) as $\hat{X}=U^{\prime}\setminus D_{\infty}$.
Define
(2.1) $X_{esc}=\{x\in\hat{X}:g^{n}(y)=x\mbox{ for some }n\in\mathbb{N},\ y\in D_{\infty}\}.$
As $D_{\infty}$ is open and $g$ is continuous, $X_{esc}$ is open (rel. $\hat{X}$): for every $x\in X_{esc}$ there exist a disk $B_{x}$ centered at $x$ and some $n(x)>0$ such that $B^{\prime}_{x}\subset D_{\infty}$, for a component $B^{\prime}_{x}$ of $g^{-n(x)}(B_{x})$. Hence, we can define the following open subset of $U$ by either of two equalities:
$D_{*}=D_{\infty}\cup\cup_{x\in X_{esc}}\cup_{n\geq 0}g^{-n}(B_{x})=D_{\infty}\cup\cup_{x\in X_{esc}}BO(x).$
Let
$X_{*}=U\setminus D_{*},$
a closed subset of $\hat{X}$. As $g(\hat{X})\subset\hat{X}$ so that $GO(x)$ is well-defined for all $x\in\hat{X}$, we have equivalently:
(2.2) $X_{*}=\{x\in\hat{X}:GO(x)\subset\hat{X}\}$.
Step II. $g^{-1}(D_{*})\subset D_{*}$ and $g^{-1}(X_{*})=X_{*}$.
Indeed, as $g^{-1}(D_{\infty})\subset D_{\infty}$, the first inclusion follows from the definition of $D_{*}$. That $X_{*}$ is completely invariant follows from (2.2).
Step III. $X_{*}$ is a full compact, i.e., $X_{*}=\hat{X}_{*}$.
[…]
(2.3) $g^{-1}(\partial\hat{X}_{*})\subset\partial\hat{X}_{*}$.
[…]
As $D_{\infty}\cap\hat{X}$$\emptyset$ while $\hat{X}_{*}\subset\hat{X}$, (2.4) implies, by the definition of $X_{*}$, that $Int(\hat{X}_{*})\subset X_{*}$. This ends the proof of Step III and claims (1)-(3) of the proposition.
Step IV. It remains to show that either $P_{g}\subset X_{*}$ or $\partial X_{*}\subset P_{g}$. So assume that $X_{*}\neq\emptyset$ and $P_{g}\nsubseteq\partial X_{*}$.
First, we claim that for every $j\geq 0$ there exists a repelling periodic point inside of $U_{-j}$ in any neighborhood of $\partial U_{-j}$. Indeed, if this is not the case for some $j$, then the alternative (B) of Proposition 2.2 must hold. As It follows now from Proposition 2.2 that since $\hat{X}\subset\cap_{i\geq 0}U_{-i}$, then $\hat{X}$ is a subset of the filled Julia set $K_{g}$ of a polynomial-like restriction of the map $g:U^{\prime}\to U$. On the other hand, as $g$ is conjugate near $K_{g}$ to a polynomial of degree $d$ which is not a power map $z^{d}$, the only completely invariant subset of $K_{g}$ is $K_{g}$ itself and then $P_{g}\subset K_{g}=X_{*}$, a contradiction.
Now we assume by a contradiction that there exists $z\in\partial X_{*}\setminus P_{g}$. Consider two cases:
(a) $z\in\partial\hat{X}$. As $\partial\hat{X}=\partial D_{\infty}$, there exists a ball $B$ centered at $z$ such that $B\subset U_{-j}$ for some $j\geq 0$ and some $x\in B\cap D_{\infty}$. In particular, the family $\mathcal{G}$ of all branches of all $g^{-n}$ are well-defined univalent functions in $B$ and don’t take values in the set $\cup_{k\geq 0}g^{-k}(P_{g})$. Hence, $\mathcal{G}$ is a normal family. On the other hand, we proved that there exists a repelling periodic point $a_{j}\in D_{\infty}\cap U_{-j}$. As $x\in B\cap D_{\infty}$, by Lemma 2.4, there exists a sequence of branches $g^{-n_{k}}\in\mathcal{G}$ such that $g^{-n_{k}}\to a_{j}$ uniformly on a compact in $B$ that contains $z\in X_{*}$. But this is a contradiction as $X_{*}$ is completely invariant and away from $a_{j}$.
[…]
3. Proof of Theorem 1.2
Assume the contrary and let $g^{p}(a)=a$ and $|\lambda|>1$ for some $p\in\mathbb{N}$, $a\in U^{\prime}$ where $\lambda=(g^{p})^{\prime}(a)$. Replacing the map $g:U^{\prime}\to U$ by $g^{p}:g^{-p}(U)\to U$ and $d$ by $d^{p}$, we can assume that $p=1$, that is, $a$ is a repelling fixed point of $g$.
Let $\psi$ be a local branch of $g^{-1}$ that fixes $a$ and is defined in a small neighborhood $Z$ of $a$ such that $\overline{\psi(Z)}\subset Z$. For each $n>0$, $\psi^{n}(Z)\subset Z$, $\psi^{n}(a)=a$ and $(\psi^{n})^{\prime}(a)=\lambda^{-n}$. One can also assume that the Konigs linearization function $S:Z\to\mathbb{C}$ of $\psi$ maps $Z$ onto a small disk centered at zero, where $S:Z\to\mathbb{C}$ is a univalent map such that $S^{\prime}(a)\neq 0$ and $S\circ\psi=\lambda^{-1}S$ on $Z$.
3.1. The case of connected $K$
Let $\pi:\mathbb{D}^{*}:=\{|w|>1\}\to\mathbb{C}\setminus K$ be the conformal isomorphism such that $\pi(\infty)=\infty$. Let $\hat{g}:=\pi^{-1}\circ g\circ\pi:\hat{U}^{\prime}\to\hat{U}$ where $\hat{U}^{\prime}=\pi^{-1}(U^{\prime}\setminus K)$, $\hat{U}=\pi^{-1}(U\setminus K)$.
Let us fix a simply connected domain $W$ in $\mathbb{D}^{*}$ such that, if $\hat{Z}=\pi^{-1}(Z)$, then $\hat{a}\in\hat{Z}\subset W$ and $\partial W\cap S^{1}$ contains a nontrivial arc $l$ of $S^{1}$ Since all branches $\hat{g}^{-n}$ are well-defined in $W$, for every $n$, the local map $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$ admits an analytic continuation along every curve in $W$ with the initial point $\hat{a}$. Therefore, every $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$. extends to a univalent map $\hat{\psi}_{n}:W\to\mathbb{C}$. Let us "exhaust" $W$ by an increasing sequence of simply connected domains $W_{i}$, $i=1,2,$..., such that $\overline{W_{i}}\subset W$, for each $i$. As $(\hat{\psi}_{n})^{\prime}(\hat{a})=(\hat{\psi}^{n})^{\prime}(\hat{a})\to 0$, by Koebe’s distortion theorem (or simply by the normality argument), for every $i$ there exists $n_{i}$ such that $\hat{\psi}_{n_{i}}(W_{i})\subset\hat{Z}$. This allows us to extend the Konigs linearization function $\hat{S}:\hat{Z}\to\mathbb{C}$ of $\hat{\psi}:\hat{Z}\to\hat{Z}$ to a holomorphic (in fact, univalent) function $\hat{S}_{i}:W_{i}\to\mathbb{C}$ according to the equation $\hat{S}_{i}(w)=\lambda^{n_{i}}\circ\hat{S}\circ\hat{\psi}_{n_{i}}(w)$. In this way, $\hat{S}$ extends to a holomorphic function $\hat{S}:W\to\mathbb{C}$. On the other hand, the above equation implies easily that $\hat{S}(w)\to\infty$ as $w\to\partial W\cap S^{1}$. This is a contradiction with the Riesz Uniqueness theorem (or more general Privalov’s theorem), e.g. [8].
3.2. Proof of Theorem 1.2 for disconnected $K$
We apply results of Appendices A-B. By Theorem A.1 and Corollary A.1, the logarithmic capacity of the completely invariant compact set $K$ is positive so Corollary B.1 applies. Let us fix an open subset of $\mathbb{C}$ as follows:
- it contains the fixed point $a$,
- it consists of finitely many, but at least two, Jordan domains $Q_{1},...,Q_{m}$, $m\geq 2$, such that $\overline{Q}_{i}$, i=1,,mi=1,...,m, are pairwise disjoint and $K\subset\cup_{i=1}^{m}Q_{i}$.
Let $\varepsilon>0$ be the distance between the set $\cup_{i=1}^{m}\partial Q_{i}$ and $K$. For this $\varepsilon$, we find a simply connected domain $W$ as in Corollary B.1, in particular, the connected set $\pi(W)\subset\mathbb{C}\setminus K$ must lie in one and only one of the domains $Q_{1},...,Q_{m}$. W.l.o.g. one can assume that
$\overline{\pi(W)}\subset Q_{1}.$
Let us joint the point $a$ and a point $q\in\partial\pi(W)\setminus K$ by a simple arc $\alpha$ which, on the one hand, lies in the range $U_{0}$ of $g$, and, on the other hand, except its end point $q$, lies outside of $\overline{\pi(W)\cup K}$ as well as outside of all others $\overline{Q_{i}}$, $2\leq i\leq m$, and also crosses the Jordan curve $\partial Q_{1}$ at a unique point. Choose a point $\hat{q}\in\partial W\cap\mathbb{D}$ such that $\pi(\hat{q})=q$ and let $\hat{\alpha}$ be the lift of the curve $\alpha$ by the covering $\pi$ that begins at $\hat{q}$. Then $\tilde{\alpha}$ is also a simple arc (in $\mathbb{D}$) with the end points $\hat{q}$ and some $\hat{a}\in\mathbb{D}$ such that $\pi(\hat{a})=a$. Note that $\hat{\alpha}\cap\overline{W}=\{\hat{q}\}$. Let us slightly "thicken" the curve $\hat{\alpha}$ to a narrow "tube" $T$ so that $\Omega:=T\cup W$ is a simply connected domain in $\mathbb{D}$ which contains the point $\hat{a}$. Then $\pi(\Omega)=\pi(T)\cup\pi(W)$ is a domain in $\mathbb{C}\setminus K$ which is disjoint with $\overline{Q_{i}}$, for all $2\leq i\leq m$. As $\alpha$ crosses $\partial Q_{1}$ at a single point, we can also assume that $Q:=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with all others $Q_{2},...,Q_{m}$.
Let $\psi:Z\to Z$ be the local inverse branch of $g$ introduced at the beginning of Sect. 3. As the neighborhood $Z$ of the fixed point $a$ can be chosen small enough and $\pi(\hat{a})=a$, we can "lift" the map $\psi:Z\to Z$ to a univalent map $\hat{\psi}:\hat{Z}\to\hat{Z}$ where $\hat{Z}$ is a neighborhood of $\hat{a}$, so that $\pi:\hat{Z}\to Z$ is a homeomorphism and $\pi\circ\hat{\psi}=\psi\circ\pi$ on $\hat{Z}$. For each $n\in\mathbb{N}$, consider the $n$-iterate $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$.
Claim 0. For each $n$, $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$ can be extended to a well-defined holomorphic function (which we call $\hat{\psi}_{n}$) in the domain $\Omega$.
Indeed, let $\hat{\beta}$ be any curve in $\Omega$ that starts at $\hat{a}$. Then $\beta:=\pi\circ\hat{\beta}$ is a curve in $\pi(\Omega)$ that starts at $a$ and lies in $U_{0}\setminus K$. As $g^{n}:U_{-n}\setminus K\to U_{0}\setminus K$ is an unbranched covering map, the local branch $\psi^{n}:Z\to Z$ of $g^{-n}$ has an analytic continuation along the curve $\beta$, and it is lifted by the unbranched covering map $\pi:\mathbb{D}\to\hat{\mathbb{C}}\setminus K$ to an analytic continuation of $\hat{\psi}^{n}$ along the curve $\hat{\beta}$. Since the latter is any curve starting at $\hat{a}$ in the simply connected domain $\Omega$, by the Monodromy theorem, the function $\hat{\psi}_{n}$ is well-defined in the whole domain $\Omega$ and $\hat{\psi}_{n}=\hat{\psi}^{n}$ in $\hat{Z}$.
Moreover,
Claim 1. For each $n$, the function $\hat{\psi}_{n}:\Omega\to\mathbb{C}$ is univalent.
Indeed, assume the contrary: $\hat{\psi}_{n}(x_{1})=\hat{\psi}_{n}(x_{2})$ for some $x_{1}\neq x_{2}$ in $\Omega$. Let $\hat{l}$ be a simple curve in $\Omega$ joining $x_{1}$ and $x_{2}$. As all branches of $g^{-n}$ are inverse branches of a well-defined unbranched covering map $g^{n}:U_{-n}\setminus K\to U_{0}\setminus K$ and since $Q=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with some other $Q_{j}$, $j\neq 1$, the following must hold:
  • $\pi(x_{1})=\pi(x_{2})$, i.e., $l:=\pi\circ\hat{l}$ is a closed curve in $\pi(\Omega)\subset Q\setminus K$,
  • an analytic continuation of some local branch of $g^{-n}$ along the closed curve $l$ maps this curve again onto some closed curve $L$,
  • there exists a lift of $L$ by $\pi$ to a curve $\hat{L}$ which begins and ends at the same point $\hat{y}:=\hat{\psi}_{n}(x_{1})=\hat{\psi}_{n}(x_{2})$, in other words the lift of $L$ is again a closed curve.
The latter condition implies that $L$ must be homotopic to a point in the domain $\hat{\mathbb{C}}\setminus K$. On the other hand, as the curve $l$ lies in a simply connected domain $Q$ which is disjoint with some $Q_{j}$, the curve $L$ must lie in one of the components of $g^{-n}(Q)$, which, in turn, are disjoint with all components of $g^{-n}(Q_{j})$. As each of the latter components contains points of $K$, the curve $L$ cannot enclose $\infty$ (in $\hat{\mathbb{C}}$), hence, $L$ is homotopic to a point in the subdomain $\mathbb{C}\setminus K$ of $\mathbb{C}$. Since $K$ is completely invariant, then $l$ is homotopic to a point in $\mathbb{C}\setminus K$ too, hence, its lift $\hat{l}$ must be a closed curve, a contradiction. This proves Claim 1.
Thus, for every $n\in\mathbb{N}$, $\hat{\psi}_{n}:\Omega\to\mathbb{D}$ is a univalent functions such that $\hat{\psi}_{n}|_{\hat{Z}}=\hat{\psi}^{n}$. Let $\hat{S}:\hat{Z}\to\mathbb{C}$ be the Konigs linearization function of $\hat{\psi}$ at its attracting fixed point $\hat{a}$, so that $\hat{S}(\hat{a})=0$, $\hat{S}^{\prime}(\hat{a})\neq 0$ and $\hat{S}(w)=\lambda\hat{S}(\hat{\psi}(w))$ for $w\in\hat{Z}$. Let $\hat{S}=S\circ\pi$ on $\hat{Z}$.
Our goal is to prove that
  1. (i)
    $\hat{S}$ extends to a holomorphic function to $\Omega$ (denoted also by $\hat{S}$), and
  2. (ii)
    $\hat{S}(w)\to\infty$ as $w\to F_{0}=\partial\Omega\cap\partial\mathbb{D}$.
As the boundary of the simply connected domain $\Omega=W\cup T\subset\mathbb{D}$ is a rectifiable curve and the length of $F_{0}=\Omega\cap\partial\mathbb{D}=\partial W\cap\partial\mathbb{D}$ is positive, (i)-(ii) would be in a contradiction with Privalov’s theorem.
Proof of (i): choose an increasing sequence $W_{i}$, $i\geq 1$, of subdomains of $W$ along with an increasing sequence of "tubes" $T_{i}$, such that, for $\Omega_{i}=T_{i}\cup W_{i}$, $\overline{\Omega_{i}}\subset\Omega$ for all $i$ and $\cup_{i}\Omega_{i}=\Omega$. As $(\hat{\psi}_{n})^{\prime}(\hat{a})=\lambda^{-n}\to 0$ as $n\to\infty$ and $\hat{\psi}_{n}(\hat{a})=\hat{a}$, by Koebe’s distortion theorem applied to maps $\psi_{n}:\Omega\to\mathbb{C}$, for any fixed $i$, $\hat{\psi}_{n}(W_{i})\to\hat{a}$ as $n\to\infty$. Choose some $n_{i}$ such that $\hat{\psi}_{n_{i}}(\Omega_{i})\subset\hat{Z}$ and then extend the function $\hat{S}:\hat{Z}\to\mathbb{C}$ to some function $\hat{S}_{i}:\Omega_{i}\to\mathbb{C}$ as follows: $\hat{S}_{i}(w)=\lambda^{n}\hat{S}(\hat{\psi}_{n_{i}}(w))$ for $w\in\Omega_{i}$. It is easy to check that $\hat{S}_{i}$ is well-defined and holomorphic in $\Omega_{i}$ and $\hat{S}_{i}=\hat{S}$ on $\hat{Z}$. Then, by the Uniqueness theorem, $\hat{S}_{i}=\hat{S}_{i+1}$ on $\Omega_{i}$. That proves (i).
Thus $\hat{S}:\Omega\to\mathbb{C}$ such that $S|_{\Omega_{i}}=S_{i}$ defines a holomorphic extension of $\hat{S}:\hat{Z}\to\mathbb{C}$.
Proof of (ii): let $w_{k}\in\Omega$, $k=1,2,$..., be any sequence that tends to $F_{0}=\partial\Omega\cap\partial\mathbb{D}$. As $F_{0}=\partial W\cap\partial\mathbb{D}$, then all but finitely many of $w_{k}$ are in $W$. Let $w_{k}\in W_{i_{k}}$ and $m_{k}:=n_{i_{k}}$ so that $\hat{\psi}_{m_{k}}(W_{i_{k}})\subset\hat{Z}$ and $\hat{S}(w_{k})=\lambda^{m_{k}}\hat{S}(\hat{\psi}_{m_{k}}(w_{k}))$. Let
[…]
in other words,
[…]
Claim 2. $e_{k}\leq m_{k}$ for all $k$ large enough.
[…]
Claim 3. $m_{k}-e_{k}\to\infty$ as $k\to\infty$.
[…]
(3.1) $g^{-(m_{k}-e_{k})}|_{loc}(z_{k})\in Z.$
By Corollary B.1(6) (applied to $X=K$), as $w_{k}\to\partial\mathbb{D}$ inside of $W$, then $z_{k}=\pi(w_{k})\to K$. But the open set $Z$ is of a positive distance away from the compact $K$ and $g^{-n}(K)=K$ for all $n$. Then (3.1) can hold only if $m_{k}-e_{k}\to+\infty$.
End of proof of (ii):
$\hat{S}(w_{k})=\lambda^{m_{k}}\hat{S}(\hat{\psi}_{m_{k}}(w_{k}))\in$
[…]
4. Proof of Theorem 1.1
Choose a Jordan domain $V_{0}\subset U$ such that $K\subset V_{0}$, therefore, $P_{g}\subset V_{0}$ as well. By Theorem 1.2, the alternative (B) of Proposition 2.2 holds, which implies that the conclusion of Theorem 1.2 holds.
5. Proof of Corollary 1.1
Let $Y$ be any non-empty completely invariant full compact set.
(2) If $P_{g}\subset Y$, then Theorem 1.1 implies that $\partial Y$ is the filled Julia set of a polynomial-like restriction to a neighborhood $P_{g}$. This implies that $Y$ must coincide with the limit set of $\{g^{-n}(z)\}_{n>0}$, where $z$ can be taken to be e.g. a critical point of $g:U^{\prime}\to U$. Hence, $Y$ is unique in this case, i.e., $K_{*}=Y$. To prove the second claim, let $x\in U$ be such that $O(x)$ is well-defined but $\omega(x)\nsubseteq\partial U$. In other words, there exist a sequence $n_{i}\to\infty$ a compact $E\subset U$ so that $g^{n_{i}}(x)\in E$, for all $i$. Choose a Jordan domain $V$ such that $P_{g}\cup E\subset V\subset\overline{V}\subset U$. As $P_{g}\subset K_{*}$, by Theorem 1.2 along with Proposition 2.2, the alternative (B) of Proposition 2.2 must hold, i.e., for some $n\in\mathbb{N}$, $V_{-n}:=g^{-n}(V_{0})$ is contained in $V$ with its closure. By (B), then $K_{*}=K_{G}:=\cap_{k\geq 0}g^{-kn}(V)$. On the other hand, write $n_{i}=l_{i}n+r_{i}$, with integers $l_{i},r_{i}$ such that $l_{i}\to\infty$ and $0\leq r_{i}<n$. Passing to a subsequence, one can assume that $r_{i}=r$ for all $i$. Then $g^{l_{i}n}(g^{r}(x))\in E\subset V$, for all $i$, hence, $g^{r}(x)\in\cap_{i\geq 0}g^{-l_{i}n}(V)\subset\cap_{k\geq 0}g^{-kn}(V)$, where the latter equality follows from the fact that the sequence of domains $\{g^{-kn}(V)\}_{k\geq 0}$ is decreasing. Thus $g^{r}(x)\in K_{*}$ so that $x\in g^{-r}(K_{*})=K_{*}$.
(3) Assume that $P_{g}\nsubseteq Y$. Let $X=P_{g}\cup Y$. Then $X$ is a forward invariant compact subset of $U$ which contains $P_{g}\cup Y$. Define $X_{*}$ as in Proposition 2.5. Then $Y\subset X_{*}$. On the other hand, by Proposition 2.5, either $P_{g}\subset X_{*}$ and then, by part (1) of the proof, $Y=X_{*}\supset P_{g}$, a contradiction, or $\partial X_{*}\subset P_{g}$. The latter implies that $K_{*}=X_{*}=\{z:GO(z)\subset\hat{P}_{g}\}\supset Y$. The conclusion follows.
Now, let $Z$ be any non-empty completely invariant compact set and let, similar to above, $X=Z\cup P_{g}$. Then $X$ is forward invariant. Define $X_{*}$ as in Proposition 2.5. By Proposition 2.5 and by the proven part, $K_{*}=X_{*}$. On the other hand, $Z\subset X_{*}$.
The final claim involving the set $R_{g}$ follows directly from Corollary 2.3 in cases (1) and (3), and from the Straightening theorem coupled with the fact that repelling cycles are dense in the polynomial Julia set in case (2).
6. Proof of Theorem 1.3
The proof follows almost literally the one of Theorem 1.1 for disconnected $K$, where very minor changed are required and only in Proposition 2.2 and its proof. First, one should restrict it only to Jordan curves $\Gamma_{0}$ which are close enough to $\partial U$. Curves $\Gamma_{n}$ turn into multi-curves. In the alternative (C), $\overline{V_{0}}$ should be a subset of a component of $V_{-n}$. In the part of the proof of Proposition 2.2 where (A) holds, the normality argument in the case $P_{g}=\emptyset$ can be easily amended; in the part where (B) holds, $V$ should be a neighborhood of $K$ consisting of a finitely many simply connected components and the Second proof of the construction of $V$ such that $g^{-1}(V)$ is compactly in $V$ should be chosen.
Appendix A Entropy of locally analytic maps
The following result is contained essentially in [5], [6]:
Theorem A.1.
(1) Let $X\subset\mathbb{C}$ be a compact subset of the plane and $f:X\to X$ be a continuous map which extends to a holomorphic map $f:W^{\prime}\to W$ of some neighborhoods $W^{\prime},W$ of $X$, such that $f^{-1}(X)=X$ and $\deg f|_{X}=N\geq 2$. Assume that $X$ is not equal to the union of cycles of critical points of $f$. Then $h_{top}(f)\geq\log N$, for the topological entropy $h_{top}(f)$ of $f:X\to X$.
(2) If, additionally, $f:W^{\prime}\to W$ is a proper map of degree $N$ where $W$ is simply connected and $W^{\prime}$ consists of a finitely many simply connected components, then the conclusion holds unless $W^{\prime}$ is (simply) connected and $X$ is a singleton $\{c\}$ where $c$ is a fixed critical point of $f$ of multiplicity $N$.
Proof.
(1) We follow closely [6] (where $X$ is the Julia set of a rational function $f$) and [5] (where $f$ is a smooth self-map of a compact smooth manifold) taking into account that our map $f$ is defined only in a small neighborhood of $X$. Shrinking $W$ if necessary one can assume that all critical points of $f:W^{\prime}\to W$ lie in $X$. Let $C$, $C_{s}$ be the sets of non periodic and periodic critical points of $f:X\to X$, respectively. (It is possible though that either $C_{s}$ or both sets $C,C_{s}$ are empty.) $C(\sigma)=\cup_{c\in C}B(c,\sigma)$ be the $\sigma$-neighborhood of $C$ and $C_{s}(\sigma)$ be a neighborhood of $C_{s}$ which is forward invariant ($f(C_{s}(\sigma))\subset C_{s}(\sigma)$) and is contained in the $\sigma$-neighborhood of $C_{s}$ (all distances are Euclidean). Note that $C(\sigma)$ is disjoint with $C_{s}(\sigma)$ for all $\sigma>0$ small enough. We prove:
($\star$) given $t\in(0,1)$, there exist $\sigma>0$ small enough and $k_{0}$ such that for every $k\geq k_{0}$ and every $y\in K$, if $y,f(y),f^{2}(y),...,f^{k}(y)\notin C_{s}(\sigma)$ then
(A.1) $\{j:0\leq j<k,f^{j}(y)\in C(\sigma)\}<tk.$
First, if $y$ is attracted by an attracting (not superattracting) or parabolic cycle, such that some $c\in C$ is also in the basin of attraction of the same cycle (the number of such cycles is at most #C\#C), then (A.1) holds because the left-hand side of (A.1) is bounded, uniformly over $y$ provided $\sigma$ is small enough. Hence, one can assume that
($\star\star$) $y$ is not attracted by a cycle that attracts also a point of $C$.
In this case, since $C,C_{s}$ are finite sets, (A.1) would follow if we show that
(A.2) $\inf\{m:z,f^{m}(z)\in B(c,\sigma)\mbox{ for some }c\in C,z\in X\}\to\infty$
as $\sigma\to 0$. To this end, let $r>0$ be such that $X_{r}:=\cup_{a\in X}B(a,r)\subset W^{\prime}$ and $M=\sup\{|f^{\prime}(z)|:z\in X_{r}\}<\infty$. Given $c\in C$, there exist $\rho>0$, $L>0$, $l\geq 2$ (the order of $c$) such that $L^{-1}|z-c|^{l-1}<|f^{\prime}(z)|<L|z-c|^{l-1}$ for all $z\in B(c,\rho)$. We prove that there exists $L_{*}$ that depends on $M$, $L$ and $l$ such that $m>L_{*}/\log(1/\sigma)$ whenever $\sigma<\min\{r/8,\rho\}$ and $z,f^{m}(x)\in B(c,\sigma)$ for some $x\in X$. It is enough to prove that
$L(5\sigma)^{l}M^{m}\geq 2\sigma.$
If we assume the contrary, then, on the one hand, all images $f^{j}(B(x,4\sigma))$, j=0,1,2,,mj=0,1,2,...,m, are well-defined, i.e. are contained in $X_{r}\subset W^{\prime}$. On the other hand, as $|x-f^{m}(x)|<2\sigma$, we obtain that $f^{m}(B(x,4\sigma))\subset B(x,4\sigma)$ which implies that $f^{mk}$ tends as $k\to\infty$ to an attracting fixed point of $f^{m}$ uniformly on compacts in $B(x,4\sigma)$. Note that $c\in B(x,4\sigma)$, therefore, $x$ and $c$ are attracted by an attracting cycle of $f$, in contradiction with ($\star\star$). This proves ($\star$).
We proceed by fixing any $\alpha\in(0,1)$. For $t=1-\alpha$, find $\sigma>0$ small enough and $k_{0}$ such that ($\star$) holds and such that the set $B(\sigma):=X\setminus C_{s}(\sigma)$ is not empty (the latter is possible by the condition on $X$). Then, for every $y\in X$ and every $k>k_{0}$,
(A.3) $\{j:0\leq j<k,f^{j}(y)\in B(\sigma)\}>\alpha k$
whenever $y,f(y),...,f^{k}(y)\notin C_{s}(\sigma)$. On the other hand, every $x\in B(\sigma)$ has a small neighborhood $U_{x}$ such that $f:U_{x}\to\mathbb{C}$ is a homeomorphism. Let $\delta>0$ be the Lebesgue number of the covering $(U_{x})$ of $B(\sigma)$. Then
(A.4) $x_{1},x_{2}\in B(\sigma),\ |x_{1}-x_{2}|<\delta\Rightarrow f(x_{1})\neq f(x_{2}).$
Now, take any $x\in B(\sigma)$. As $C_{s}(\sigma)$ is forward invariant, $B(\sigma)$ is backward invariant. It follows from the condition on $X$ that we can choose $x$ not to be in the forward orbit of any critical point of $f$.
We are in a position to repeat literally the end of the proof of [5] (where $B=B(\sigma)$).
A set $P\subset X$ is called $\delta$-separated if $|z_{1}-z_{2}|>\delta$ for all distinct $z_{1},z_{2}\in P$ and $P$ is $(k,\delta)$ separated if $\max\{|f^{j}(z_{1})-f^{j}(z_{2})|>\delta,j=0,...,k-1\}$ for all distinct points $z_{1},z_{2}\in P$ (in particular, any singleton is $(k,\delta)$-separated). Consider the backward orbit $BO(x)=\cup_{k=0}^{\infty}f^{-k}(x)$ of $x$. Then $f^{-1}(y)=N$ for all $y\in BO(x)$. For every $y\in BO(x)$ there exists a $\delta$-separated set $P(y)\subset f^{-1}(y)$ such that either $P(y)=\{z\}$ for some $z\notin B(\sigma)$ or $P(y)\subset B(\sigma)$ and #P(y)=N\#P(y)=N. Fix $k>k_{0}$ and define by induction:
$Q_{0}=\{x\};\ \ Q_{j+1}=\cup_{y\in Q_{j}}P(y),\ \ j=0,1,...,k-1.$
Then $Q_{k}$ is $(k,\delta)$-separated (induction in j=0,,kj=0,...,k). Let us show that $Q_{k}\geq N^{\alpha k}$. For $y\in Q_{j}$, let $Gen(y,j)=\#\{i:0\leq i<j,f^{i}(y)\in B(\sigma)\}$. Let $m=[\alpha k]+1$ (where $[T]$ is the entire part of $T>0$). Then, by (A.3), there are $N^{m}$ pairs $(y,j)$ such that $y\in B(\sigma)\cap Q_{j}$ and $Gen(y,j)=m$. For each such a pair there exists $z\in Q_{k}$ for which $f^{k-j}(z)=y$ and such points are distinct for distinct pairs. Therefore, $Q_{k}\geq N^{m}\geq N^{\alpha k}$. We have shown that for every $\alpha\in(0,1)$ there exist $k_{0}$ and $\delta>0$ such that the maximal cardinality $r_{k}(f,\delta)$ of a $(k,\delta)$-separated subset of $X$ is bigger than or equal to $N^{\alpha k}$, for all $k>k_{0}$. Thus $h_{top}(f)=\lim_{\delta\to 0}\limsup_{k\to\infty}\frac{1}{k}\log r_{k}(f,\delta)\geq\alpha\log N$, for all $\alpha\in(0,1)$. We are done.
(2) Suppose that $W$ is simply connected and $X\subset W^{\prime}\cap W$ consists of cycles of critical points of $f$. Then $f^{-1}(x)=1$ for all $x\in X$ which, along with the Riemann-Hurwitz formula for the map $f:U\to W$ where $U$ is a component of $W^{\prime}$ yield that $W^{\prime}$ consists of a single component and $X$ is just a point (which is then a fixed critical point of $f:W^{\prime}\to W$ of the order $d$).
Corollary A.1.
Under the conditions of Theorem A.1, the logarithmic capacity of the compact set $X$ is positive.
Proof.
By the Variational Principle, $\log N\leq h_{top}(f)=\sup h_{\mu}(f)$, where supremum is taken over the set $M(f)$ of all $f$-invariant Borel probability measures on $X$ and $h_{\mu}(f)$ is the entropy of $f$ w.r.t. $\mu$, see e.g. [7]. In particular, there exists an ergodic $\mu\in M(f)$ with $h_{\mu}(f)>0$. This implies that the Lyapunov exponent $\chi_{\mu}(f):=\int\log|f^{\prime}|d\mu>0$. Therefore, $HD(\mu)=h_{\mu}(f)/\chi_{\mu}(f)>0$, where $HD(\mu)$ is the infimum of Hausdorff dimensions over all Borel $Y\subset X$ such that $\mu(Y)=1$. Hence, Hausdorff dimension of $X$ is positive. In turn, this implies (see e.g. [8]) that the logarithmic capacity of $X$ is positive as well. ∎
Remark A.2.
By [7], Theorem 11.6.1, any such ergodic $\mu\in M(f)$ is approximated by measures supported on Cantor expanding repellers of $f$. In particular, there exist a lot of periodic orbits in any neighborhood of $X$. However, in general they are off $X$.
Appendix B Uniformization
Theorem B.1.
(see e.g. [8]) Let $X$ be a full compact set in $\mathbb{C}$ that contains at least $3$ points. Let $\pi:\mathbb{D}=\{|w|<1\}\to\hat{C}\setminus X$ be a universal holomorphic covering map and $\Gamma=\{\gamma\}$ be the corresponding Fuchsian group of Mobius transformations of $\mathbb{D}$ without elliptic elements such that $\mathbb{C}\setminus X\simeq\mathbb{D}/\Gamma$. Normalize $\pi$ such that $\pi(0)=\infty$. The following conditions are equivalent:
  1. (1)
    logarithmic capacity $c(X)$ of $X$ is positive,
  2. (2)
    $\Gamma$ is of convergence type, that is
    $\sum_{\gamma\in\Gamma}(1-|\gamma(w)|)<\infty$
    for some (hence, any) $w\in\mathbb{D}$. In this case, the infinite Blaschke product
    $B(w)=\Pi_{\gamma\in\Gamma}\frac{|\gamma(0)|}{\gamma(0)}\frac{w-\gamma(0)}{1-\overline{\gamma(0)}w}$
    is a holomorphic function in $\mathbb{D}$ such that $|B\circ\gamma(w)|=|B(w)|$ for all $w\in\mathbb{D}$, $\gamma\in\Gamma$, $|B(w)|<1$ for all $w\in\mathbb{D}$ and, for almost every $w_{0}\in\partial\mathbb{D}$, the angular limit $\lim|B(w)|=1$ as $w\to w_{0}$ exists.
    Moreover, projecting the function $\log|B|$ to $\mathbb{C}\setminus X$ we obtain Green’s function $G_{X}$ of $\mathbb{C}\setminus X$, that is,
    $G_{X}(z)=-\log|B(w)|,$
    for some (hence, every) $w\in\mathbb{D}\setminus\{\gamma(0):\gamma\in\Gamma\}$, such that $z=\pi(w)$. Then $G_{X}$ is well-defined everywhere positive harmonic function in $\mathbb{C}\setminus X$, $G(z)-\log|z|\to-\log(c(X))$ as $z\to\infty$ and, for almost every $w_{0}\in\partial\mathbb{D}$, the following angular limit exists:
    (B.1) $\lim_{w\to w_{0}}G_{X}(\pi(w))=0.$
For a Borel subset $F\subset\partial\mathbb{D}$, let $|F|$ be its Lebesgue measure (length). We have:
Corollary B.1.
Let $X$ be a compact with positive logarithmic capacity. Then for every $\varepsilon>0$ there exists a domain $W\subset\mathbb{D}$ as follows:
  1. (1)
    $W$ is simply connected,
  2. (2)
    $\partial W$ is a rectifiable Jordan curve,
  3. (3)
    there exists a point $w_{0}\in\partial W$ such that $|B(w)|>|B(w_{0})|$ for all $w\in W$ and $\pi(w)=\pi(w_{0})$ for finitely many $w\in\partial W$,
  4. (4)
    $|\partial W\cap\partial\mathbb{D}|>0$,
  5. (5)
    $\pi(W)$ lies in the $\varepsilon$-neighborhood of $X$,
  6. (6)
    $\pi(w)\to X$ uniformly as $w\to\partial\mathbb{D}$ inside of $W$.
Proof.
Let $X_{\varepsilon}=\{z\in\mathbb{C}\setminus X:d(z,X)<\varepsilon\}$. Define
$l_{\varepsilon}=\inf\{G_{X}(z):z\in\mathbb{C}\setminus X_{\varepsilon}\}.$
Then $l_{\varepsilon}>0$ and if $G_{X}(z)<l_{\varepsilon}$ for some $z\in\mathbb{C}\setminus X$, then $z\in X_{\varepsilon}$.
Let $b=\exp(-l_{\varepsilon})$. Then $|B(w)|>b$ implies $\pi(w)\in X_{\varepsilon}$.
Now we repeat partly a construction as in the proof of Privalov’s theorem, e.g. [8]. Fix an angle $\beta\in(0,\pi/2)$ and given $a=\exp(i\theta)\in\partial\mathbb{D}$ and $n\in\mathbb{N}$ define
$H_{n}(a)=\inf\{|B(w)|:1-1/n<|w|<1,|\arg(w-a)|<\beta\}.$
By Theorem B.1,
$\lim_{n\to\infty}H_{n}(a)=1$
for a.e. $a\in\partial\mathbb{D}$. Hence, by Egorov’s theorem, there exists a closed subset $F\subset\partial\mathbb{D}$ of positive measure such that $H_{n}(a)\to 1$ as $n\to\infty$ uniformly in $a\in F$. In particular, for some $n_{0}\in\mathbb{N}$,
$|B(w)|>b$
whenever
$w\in W_{*}=\{1-1/n_{0}<|w|<1,\arg(w-a)<\beta,a\in F\}.$
One can assume that $F$ lies in an arc of $\partial\mathbb{D}$ of small enough length and $n_{0}$ is big enough so that the part of the boundary of $W_{*}$ which lies on the circle $|w|=1-1/n_{0}$ is a proper part of this circle. Note that $\partial W_{*}\cap\partial\mathbb{D}=F$. Hence, there exists a connected component $W$ of the (open) set $W_{*}$ (note that $W=W_{*}$ if $W_{*}$ is connected itself) and a closed subset $F_{0}$ of $F$ of positive measure as follows:
$W=\{1-1/n_{0}<|w|<1,|\arg(w-a)|<\beta,a\in F_{0}\},$
$|B(w)|>b$ for every $w\in W$ and $|B(w)|\to 1$ as $w\to F_{0}$ uniformly in $W$. Then properties (4)-(6) hold for $W$. For the property (3), let $w_{0}\in\overline{W}$ be so that $|B(w_{0})|=\inf\{|B(w)|:w\in\overline{W}\setminus\partial\mathbb{D}\}$. Then $w_{0}\in\partial W$ and $|B(w)|>|B(w_{0})|$, by the Minimum principle and there could be at most finitely many $w_{n}\in\partial W$ with $\pi(w_{n})=\pi(w_{0})$ as otherwise $w_{n}\to\partial W\cap\partial\mathbb{D}$ which is impossible as then $|B(w_{0})|=|B(w_{n})|\to 1$. Let’s show that (1)-(2) hold too. By the construction, $\partial W\cap\{|w|=1-1/n_{0}\}$ is a proper arc $\{(1-1/n_{0})\exp{it}:t_{-}\leq t\leq t_{+}\}$ of the circle $|w|=1-1/n_{0}$. Then the closure of $W$ is a subset of $R:=\{w=r\exp(it):1-1/n_{0}\leq r\leq 1,t_{-}\leq t\leq t_{+}\}$. Therefore, there exists a neighborhood $R^{+}$ of $R$ and a diffeomorphic in $R^{+}$ branch $\bf Log$} of the map $w\mapsto i\overline{\log(w)}$ such that (slightly abusing notation and replacing $W$ by ${\bf Log}(W)$) one can assume that $F_{0}$ is a compact subset of $\mathbb{R}$ and
$W=\{z=x+iy:0<y<h,|\arg(z-x_{0})|<\pi/2-\beta\mbox{ for some }x_{0}\in F_{0}\},$
where $h=-\log(1-1/n_{0})$. Then it is easy to see that the boundary of $W$ is connected, which implies that $W$ is simply connected. Now, if $I=[x_{1},x_{2}]\subset\mathbb{R}$ is the smallest closed interval containing $F_{0}$ and $I\setminus F_{0}=\cup_{j=1}^{\infty}I_{j}$ where $I_{j}$ are pairwise disjoint open intervals, then the boundary of $W$ consists of: the "upper" boundary interval $\{x+ih:x_{1}-h\tan(\beta)\leq x\leq x_{2}+h\tan(\beta)\}$, "left" and "right" intervals $x_{1}-h\tan(\beta)+ih,x_{1}]$ and $x_{2},x_{2}+h\tan(\beta)+ih]$, the compact set $F_{0}\subset\mathbb{R}$ and a "saw"-like curve consisting of a sequence of two sides of equilateral triangles with bases $I_{j}$ whose total length is bounded by $\sum_{j}|I_{j}|/\sin(\beta)=(|I|-|F_{0}|)/\sin(\beta)$. Thus $\partial W$ is a rectifiable curve.