### v2

# Polynomial-like dynamics of analytic maps

Abstract.
The theory of polynomial-like maps is of fundamental importance in holomorphic dynamics. We study dynamical properties of a larger class of maps. Our main result is that, under some natural assumptions, a map of this class has a completely invariant compact set if and only if this set is the filled Julia set of a polynomial-like restriction of the map. We also generalize this result to include maps with non-connected domains of definition.
1. Introduction and main results
Polynomial-like maps were introduced by Douady and Hubbard [4] to explain, in particular, the fractal nature of non-linear phenomena in holomorphic dynamics. In this paper proper holomorphic map which are not polynomial-like are considered and we are interested in conditions that guarantee a polynomial-like (PL) restriction of such a map. We prove in Theorem 1.1 that, in the class of BI maps, see Definition 1.1, a PL restriction exists if and only if the map has a completely invariant compact set $K$ (under some natural restrictions). This gives a rather complete description of the dynamics of BI maps on such a compact and its neighborhood. It would be interesting to study dynamical properties of locally analytic self-maps of a compact set of degree at least two for other classes of maps.
Let
(1.1) $g:U_{-1}\to U_{0}$
be a proper holomorphic map of some degree $d\geq 2$, where $U_{-1},U_{0}$ are Jordan domains. We study such a map under an extra assumption that it is backward invariant, i.e., all further (following the first one $U_{-1}=g^{-1}(U_{0})$) pullbacks $g^{-i}(U_{0})$, $i=2,3,$..., of $U_{0}$ also exist (BI map, see Definition 1.1).
If $\overline{U_{-1}}\subset U_{0}$, the map (1.1) is polynomial-like and is always BI because all pullbacks exists (see Section 1.1 and Example 1.3 there for details). On the other hand, there are two interesting classes of BI maps which are not, in general, PL maps, see Example 2.1 for the first class and Examples 2.2-2.3 for the second one. If, e.g., in (1.1), $U_{-1}\subset U_{0}$, but $\overline{U_{-1}}\cap\partial U_{0}\neq\emptyset$, and the postcritical set is a subset of $U_{-1}$, the map is BI but not PL. On the other hand, another class of BI maps appears naturally in connection with the renormalization. Namely, a BI map can be associated to any simple renormalization of a quadratic polynomial, see Example 2.2. If, moreover, a quadratic polynomial is infinitely renormalizable and robust but with no unbranched complex bounds, in the sense of McMullen [9], a rescaling subsequence of BI maps along renormalizations converges to a limit BI map, see Example 2.3. In fact, the latter examples inspired the present work.
The set $C_{g}:=\{z\in U_{-1}:g^{\prime}(z)=0\}$ of critical points of the map (1.1) is not empty as $U_{-1},U_{0}$ are simply connected and the degree $d\geq 2$. Let
$P_{g}=\overline{\cup_{n>0}g^{n}(C_{g})}$
be the postcritical set of this map whenever it is well-defined, i.e., assuming $g^{n}(C_{g})\subset U_{-1}$ for all $n\geq 0$.
Definition 1.1.
The map (1.1) is called backward invariant (BI map) if there exist a sequence
$U_{0},U_{-1},...,U_{-n},$...
of Jordan domains and a holomorphic extension $g:\cup_{i=1}^{\infty}U_{-i}\to\cup_{i=0}^{\infty}U_{-i}$ of $g:U_{-1}\to U_{0}$ such that the following hold, for every $i=0,1,$...:
  • $g:U_{-i-1}\to U_{-i}$ is a proper holomorphic map of degree $d$,
  • $P_{g}$ is well-defined and $P_{g}\subset U_{-i}$. Moreover, we assume that if $P_{g}$ is a singleton $\{c\}$ then $g^{-1}(c)\neq\{c\}$.
In the sequel, slightly abusing notation,
(1.2) $g:U^{\prime}\to U,\mbox{ where }U^{\prime}=\cup_{i=1}^{\infty}U_{-i},U=\cup_{i=0}^{\infty}U_{-i},$
always refers to the BI map (unless stated otherwise).
We call a BI map $g:U^{\prime}\to U$ trivial if there is an attracting fixed point $a\in\cap_{i\geq 0}U_{-i}$ such that $g^{n}(z)\to a$ for all $z\in U$.
For every $i\geq 0$ and $n\geq 1$, the map $g^{n}:U_{-i-n}\to U_{-i}$ is a proper map of degree $d^{n}$. By $g_{i}^{-n}:U_{-i}\to U_{-i-n}$ we denote the inverse (multi-valued) map. E.g., $g_{0}^{-n}(U_{0})=U_{-n}$.
Remark 1.2.
We do not assume that $g:U^{\prime}\to U$ is a proper map (of degree $d$). But if it is so (as in Example 2.1), then writing $g^{-n}(x)$ (i.e. dropping $i$ in $g_{i}^{-n}(x)$) is not ambiguous and this would allow us to prove a bit more, see Remark 1.6 and the Complement to Corollary 1.5. However, there is no reason to believe that $g:U^{\prime}\to U$ is a proper map in Examples 2.2-2.3.
The following maps (i)-(iii) which are associated to the BI map (1.2) are BI by themselves: (i) the restricted map $g:\cup_{n=1}^{\infty}g_{0}^{-n}(\Omega)\to\cup_{n=0}^{\infty}g_{0}^{-n}(\Omega)$, for every Jordan domain $\Omega$ such that $P_{g}\subset\Omega\subset U_{0}$, (ii) the shifted map $g:\cup_{i=i_{0}+1}^{\infty}U_{-i}\to\cup_{i=i_{0}}^{\infty}U_{-i}$, for every $i_{0}>0$, and (iii) the iterated map $g^{n}:\cup_{i=1}^{\infty}U_{-ni}\to\cup_{i=0}^{\infty}U_{-ni}$, for every $n>1$. All these maps share the same postcritical set $P_{g}$.
1.1. Polynomial-like maps
Recall that a polynomial-like (PL) map is a triple $(V^{\prime},V,g)$ where $V^{\prime},V$ are topological disks in the plane such that $\overline{V^{\prime}}\subset V$ and $g:V^{\prime}\to V$ is a proper holomorphic map of some degree $d\geq 2$ [4]. The non-escaping set $K_{g}:=\{z:g^{n}(z)\in U_{-1},n=0,1,2,$...\} is called the filled Julia set of the PL map $g$.
It is easy to see that $K_{g}$ is a compact set, it is full, i.e., its complement $\mathbb{C}\setminus K_{g}$ is connected, and is completely invariant, i.e., $g^{-1}(K_{g})=K_{g}$.
The compact $K_{g}$ is connected if and only if $C_{g}\subset K_{g}$ where $C_{g}=\{z\in V^{\prime}:g^{\prime}(z)=0\}$. Equivalently, $P_{g}\subset V$.
By the Straightening Theorem [4], a PL map $g:V^{\prime}\to V$ is hybrid equivalent (i.e., conjugate in some neighborhood of $K_{g}$ by a quasiconformal homeomorphism which is conformal a.e. on $K_{g}$) to a polynomial of degree $d$. Since the polynomial dynamics is well-understood, see e.g., [2], this gives a complete description of the dynamics on $K_{g}$. In particular, the set of repelling periodic points of the map $g:K_{g}\to K_{g}$ is dense in $\partial K_{g}$, for every $n\in\mathbb{N}$, the map $g^{n}:K_{g}\to K_{g}$ has $d^{n}$ fixed points (counting multiplicity), every component of the interior of $K_{g}$ is (pre)periodic [16], etc. Moreover, the set of limit points of $\{g^{-n}(x)\}_{n\geq 0}$ coincide with $\partial K_{g}$, for all $x\in K_{g}$ with the only exception if $g^{-1}(x)=\{x\}$ (in the latter case, $C_{g}=\{x\}$ and is completely invariant).
[…]
Every PL map is, essentially, a BI map:
Example 1.3.
Given a PL map $g:V^{\prime}\to V$ with $P_{g}\subset V^{\prime}$, choose any Jordan domain $U_{0}\subset V$ which is slightly smaller than $V$ (if $V$ is Jordan itself, one can take $U_{0}=V$). Then $U_{-i}:=g^{-i}(U_{0})$, $i=0,1,2,$..., is a well-defined (decreasing in this case) sequence of Jordan domains such that $P_{g}\subset U_{-i}$ and $g:U_{-i-1}\to U_{-i}$ is proper degree $d$, for each $i$. Assume additionally that $C_{g}$ is not a completely invariant singleton (equivalently, $g:V^{\prime}\to V$ is not hybrid equivalent to the power map $z\mapsto z^{d}$). Then $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ is a BI map.
See Section 2 for examples of BI maps which are not polynomial-like.
Example 1.3 makes natural the following
Definition 1.4.
We say that a BI map $g:U^{\prime}\to U$ has a polynomial-like restriction if there exists a topological disk $V\subset U_{0}$ as follows: if $V^{\prime}:=g_{0}^{-1}(V)$ then $P_{g}\subset V^{\prime}\subset\overline{V^{\prime}}\subset V$ and $g:V^{\prime}\to V$ is a polynomial-like map of degree $d$.
In this case, the filled Julia set $K$ of the PL map $g:V^{\prime}\to V$ is a full completely invariant compact containing $P_{g}$.
1.2. Main result
Example 1.3 provides us a necessary condition for a BI map to have a PL restriction: the existence of a full completely invariant compact set that contains all critical points. We prove that this necessary condition is also sufficient:
Theorem 1.1.
Suppose that $g:U_{-1}\to U_{0}$ is a proper holomorphic map of some degree $d\geq 2$ which extends to a BI map $g:U_{-i-1}\to U_{-i}$, $i=1,2,$.... Let $K\subset\cap_{i\geq 0}U_{-i}$ be a compact set as follows:
  1. (t1)
    $K$ is completely invariant,
  2. (t2)
    $K$ is a full compact and $P_{g}\subset K$.
Then $g$ has a polynomial-like restriction and $K$ is its filled Julia set. In particular, $K$ is a continuum,
In [5] we apply Theorem 1.1 to study holomorphic/meromorphic invariant line fields of BI maps. In the present paper we derive the following corollary and its complement.
Given a BI map $g:U^{\prime}\to U$ we define, for each $n\in\mathbb{N}$, the set
(1.3) $R_{n}=\{z|z\in U_{-n}\cap U_{0},g^{n}(z)=z,|(g^{n})^{\prime}(z)|>1\}.$
Note that $R_{n}$ is finite in any compact subset of $U_{0}$, by the Uniqueness theorem. Let
(1.4) $R^{\prime}=\{z|\exists z_{i}\to z,z_{i}\in R_{n_{i}},n_{i}\to\infty\}.$
Corollary 1.5.
For a non-trivial BI map $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$, the following alternative holds:
(a) it does not have a completely invariant full compact $K$ such that $P_{g}\subset K\subset\cap_{i\geq 0}U_{-i}$, and the set $R^{\prime}$ contains a (non trivial) continuum which has common points with $U_{0}$ as well as with its boundary $\partial U_{0}$.
(b) $g$ admits such $K$ and there exists a neighborhood of $W$ of $K$, such that $g:g_{0}^{-1}(W)\to W$ is a PL restriction and $K$ its filled Julia set. Besides,
  1. (1)
    $R^{\prime}\cap K=\partial K$,
  2. (2)
    $R_{n}\cap(W\setminus K)=\emptyset$, for every $n\in\mathbb{N}$,
  3. (3)
    $R^{\prime}\setminus K\subset\partial U_{0}$.
Remark 1.6.
Assume that a BI map $g:U^{\prime}\to U$ is a degree $d$ proper map. Then $\cup_{n\geq 1}R_{n}\subset R^{\prime}$, in particular, in case (b), $g$ has no repelling periodic orbits outside of $K$. Indeed, $g:U_{-i-1}\to U_{-i}$ has degree $d$ and if also $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ has degree $d$ then $g^{-1}(U_{-i})=U_{-i-1}$. This implies that $g_{0}^{-n}(x)=g^{-n}(x)$ for all $x\in U_{0}$ and all $n\in\mathbb{N}$. If now $z\in U_{-n}\cap U_{0}$ is such that $g^{n}(z)=z$, then $g^{ni}(z)=z$, hence, $z\in U_{-ni}$. Therefore, letting $z=z_{i}$ and $n_{i}=ni$ in (1.4) we see that $z\in R^{\prime}$.
This remark is relevant for, e.g., a class of BI maps of Example 2.1. Let us state this explicitly as the following
Complement to Corollary 1.5. Let $V^{\prime},V$ be simply connected domains in $\mathbb{C}$ such that $^{\prime}\subset V, VV$$^{\prime}\neq V$ and $g:V^{\prime}\to V$ is a proper holomorphic map of some degree $d\geq 2$ and the postcritical set $P_{g}\subset V^{\prime}$. Let $R^{\prime}$ be the closure of all repelling cycles of $g:V^{\prime}\to V$. Then the alternative (a)-(b) holds. Moreover, in case (b) of the alternative, $R^{\prime}\subset K$, and for any $x\in V^{\prime}\setminus K$, either $g^{n}(x)\in V\setminus V^{\prime}$ for some $n\geq 0$, or the forward orbit $O(x)$ of $x$ is well-defined but
$\omega(x)\subset\partial V,$
where $\omega(x)$ is the limit set of $O(x)$.
Next remark is, essentially, a folklore:
Remark 1.7.
For a proper map (1.1), conditions (t1)-(t2) along are not enough for the conclusion of Theorem 1.1 to hold. Indeed, let’s assume, along with (t1)-(t2), that $g:U_{-1}\to U_{0}$ has an attracting or parabolic cycle such that its basin of attraction intersects $U_{-1}\setminus K$. Then it is easy to see that the conclusion of Theorem 1.1 breaks down. Nevertheless, if, additionally, to (t1)-(t2), $K$ is connected and attracting/parabolic cycles as above do not exist, then the conclusion of Theorem 1.1 does hold (see [1] for details, see also Case (B) in the proof of Proposition 3.2 below). The proof of this is heavily based on the fact that $K$ is a connected full compact set. This allows us to conjugate the map $g:U_{-1}\setminus K\to U_{0}\setminus K$ (by uniformazing $\mathbb{C}\setminus K$) to a real analytic covering map on an outer neighborhood of the unit circle $S^{1}$ with no critical points, extend it through $S^{1}$ by the reflection principle and then apply a non-trivial Mane’s theorem which says that a $C^{2}$ covering map of the circle without critical points is expanding away from its attracting and parabolic cycles.
The proof of Theorem 1.1 is a mixture of dynamical and geometric arguments. The main step constitutes the following
Theorem 1.2.
Under the conditions of Theorem 1.1, the set $R^{\prime}$ (see 1.3-1.4) has no points in $U_{0}\setminus K$.
The proof of Theorem 1.2 in the case of connected $K$ is rather simple because we can still use the conjugation of the map $g:U_{-1}\setminus K\to U_{0}\setminus K$ to a circle map (although the proof is more elementary than the one outlined in the above Remark 1.7 as Mane’s theorem is not needed), see Sect. 4.1. Most part of the paper occupies the proof of Theorem 1.2 in the case of disconnected $K$ along with Appendices A-C where we collect auxiliary results which are used in the proof.
1.3. Generalized BI maps
Note that a priori the completely invariant compact $K$ of Theorem 1.1 does not have to be connected although a posteriori it does.
Let us mention the following generalization of Theorem 1.1 where $K$ is never connected. Let, in Definition 1.1, $U_{0}$ be still a Jordan domain while $U_{-1}$ be an open set consisting of finitely many Jordan components $U_{-1}^{k}$, k=1,,mk=1,...,m, for some $m\geq 2$. Let $g:U_{-1}\to U_{0}$ be a proper map of degree $d\geq 2$ meaning that each $g:U_{-1}^{k}\to U_{0}$ is a proper map of some degree $d_{k}\geq 1$ such that $d=\sum_{k}d_{k}$. We say that $g$ extends to a generalized BI map if there exist a sequence $U_{-1},...,U_{-i},$... of bounded open sets each of which consists of finitely many Jordan components, and a holomorphic extension $g:\cup_{i=1}^{\infty}U_{-i}\to\cup_{i=0}^{\infty}U_{-i}$ of $g:U_{-1}\to U_{0}$ such that, for every $i=0,1,$...: (1) $g(U_{-i-1})=U_{-i}$ and, for every component $U$ of $U_{-i}$, $g:U_{-i-1}\to U$ is a proper holomorphic map of degree $d$, (2) either $g$ has no critical points or $P_{g}$ is well-defined, contains at least two points and $P_{g}\subset U_{-i}$.
Recall that a generalized PL map [7] $G:V^{\prime}\to V$ is a holomorphic map between open sets $V^{\prime},V$, where $V$ is simply connected, $\overline{V^{\prime}}\subset V$, and $V^{\prime}$ consists of finitely many simply connected components $V_{k}^{\prime}$, k=1,,Mk=1,...,M, with pairwise disjoint closures, such that each restriction $G:V_{k}^{\prime}\to V$ is a proper map of some degree $D_{k}\geq 1$. Its set of non-escaping points (filled Julia set) $K_{G}=\{z:G^{n}(z)\in V^{\prime},n=0,1,$...\} is a completely invariant full compact. (Note that if $V^{\prime}$ is connected, $G:V^{\prime}\to V$ is the classical (Douady-Hubbard) PL map.) The Straightening theorem also holds: $G:V^{\prime}\to V$ is conjugated, in a neighborhood of $K_{G}$, to a polynomial of degree $D=\sum_{k=1}^{M}D_{k}$ (see e.g. [6] for a proof of a similar claim). If $M\geq 2$, then $K_{G}$ is never connected.
Theorem 1.3.
Suppose that $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ is a generalized BI map. Let $K\subset\cap_{i\geq 0}U_{-i}$ be a full completely invariant compact set such that either $g$ has no critical points or $P_{g}\subset K$. Then there exist a simply connected domain $V_{0}$, $\overline{V_{0}}\subset U_{0}$, and $m\geq 1$ such that $\overline{V_{-m}}\subset V_{0}$, where $V_{-m}=g_{0}^{-m}(V_{0})$, and $K$ is the filled Julia set of the generalized PL map $g^{m}:U_{-m}\to U_{0}$ of degree $d^{m}$. Furthermore, there exists a neighborhood $V\subset U_{0}$ of $K$ consisting of finitely many pairwise disjoint with their closures simply connected domains such that $\overline{g_{0}^{-1}(V)}\subset V$ and $g:g^{-1}(V)\to V$ is a proper map of degree $d$.
1.4. Description of the paper
Section 2 contains examples of BI maps which are not polynomial-like.
In Section 3 some general properties of BI maps are collected. Primarily, we prove in Proposition 3.2 that any BI map which is not trivial (i.e., not a basin of attraction of an attracting fixed point) either has a lot of repelling cycles close to the boundary and does not admit a PL restriction, or has a PL restriction, with filled Julia set containing almost all repelling cycles.
Theorem 1.2 is proved in Section 4. The proof in the case of disconnected compact $K$ is based on general facts and their corollaries from Appendices A-C.
Short Sections 5-6-7 contain proofs of Theorem 1.1 (as a simple consequence of Theorem 1.2), Corollary 1.5 along with its Complement, and Theorem 1.3, respectively.
Appendix A. We prove following closely [8], [13] that a locally analytic self-map of a compact set $K$ of degree $d\geq 2$ has topological entropy $\geq\log d$. Coupled with the Variational Principle and the formula for the Hausdorff dimension of invariant measures, this implies that the logarithmic capacity of the completely invariant compact set is positive.
Appendix B. Given a (disconnected) compact $X\subset\mathbb{C}$ of positive logarithmic capacity and a universal covering $\pi:\mathbb{D}\to\mathbb{C}\setminus X$, we recall a classical result about the existence and properties of an infinite Blaschke product $B:\mathbb{D}\to\mathbb{D}$ such that its projection is Green’s function of $\mathbb{D}\setminus X$. In Corollary B.2 we use Lusin-Privalov’s construction to build a nice simply connected domain $W\subset\mathbb{D}$, in particular, the length of $\partial W\cap\partial\mathbb{D}$ is positive, $\pi(W)$ is contained in a given neighborhood of $X$ and $\pi(w)\to X$ as $w\to\partial W\cap\partial\mathbb{D}$ in $W$.
Appendix C. We consider a univalent map $f:\mathbb{D}\to\mathbb{D}$ and show, based on an inequality proved in [10], [12], [15] that, if $E,f(E)\subset\partial\mathbb{D}$, then the length of $E$ is small provided $|f^{\prime}(0)|$ is small and $f(0)$ is away from the unit circle.
Acknowledgments. I would like to thank Feliks Przytycki for several fruitful discussions concerning Theorem A.1, and Misha Sodin, Sebastian van Strien and Vladlen Timorin for helpful comments and references.
2. Examples of BI maps
The next example generalises the previous Example 1.3. We obtain BI maps by relaxing some conditions in the definition of polynomial-like maps and in the choice of $U_{0}$.
Example 2.1.
Let $V^{\prime},V$ be simply connected (perhaps, unbounded) domains in $\mathbb{C}$ such that $^{\prime}\subset V, VV$$^{\prime}\neq V$ and $g:V^{\prime}\to V$ is a proper holomorphic map of some degree $d\geq 2$. We don’t assume that $V^{\prime}$ is compactly contained in $V$. Such a map must have a non-empty set of critical points $C_{g}=\{x:g^{\prime}(x)=0\}$. Assume that the postcritical set $P_{g}=\cup_{n>0}g^{n}(C_{g})$ is well-defined and is a compact subset of $V^{\prime}$, and if $P_{g}$ is a singleton $\{c\}$ then $g^{-1}(c)\neq\{c\}$. Let $U_{0}\subset V$ be an arbitrary Jordan domain such that $P_{g}\subset U_{0}$. Define $U_{-i}=g^{-i}(U)$, $n=0,1,2,$.... Then $g:\cup_{i\geq 1}U_{-i}\to\cup_{i\geq 0}U_{-i}$ is a BI map.
Sub-example: if $V^{\prime},V$ above are bounded and $\overline{V^{\prime}}\subset V$ (i.e., $g:V^{\prime}\to V$ is PL), then it is not difficult to see that the obtained BI map has a PL restriction if and only if $K_{g}\subset U_{0}$, where $K_{g}$ is the filled Julia set of the PL map $g:V^{\prime}\to V$ (the restriction is then $g:g^{-n}(V^{\prime})\to g^{-n}(V)$, for some $n>0$ such that $g^{-n}(V)\subset U_{0}$).
The following examples of BI maps are related to renormalizable quadratic polynomials.
Example 2.2.
Let $f(z)=z^{2}+c$ be a quadratic polynomial. Assume that it admits a simple renormalization of a period $n>1$ (see [9] for definitions). Let $J_{n}$ be the filled Julia set of a PL restriction of $g^{n}$ with a single critical point at 0, and let $P_{n}$ be the postcritical set of this PL map. Note that $P_{n}=J_{n}\cap P_{f}$. Let $\gamma_{n}$ be a Jordan curve in $M:=\hat{\mathbb{C}}\setminus P_{f}$ which is homotopic in $M$ to a loop that separates $J_{n}$ from $P_{f}\setminus J_{n}$. Let $U_{0}$ be a simply connected domain bounded by $\gamma_{n}$. Pulling back $U_{0}$ along the renormalization as in [9], we obtain a domain $U_{-1}\supset P_{n}$ and degree two proper restriction map $f^{n}:U_{-1}\to U_{0}$ with the postcritical set $P_{n}$. Similarly, we pull back $U_{-1}$ getting $U_{-2}$, and so on. As a result, we obtain a BI map $f^{n}:U_{-i-1}\to U_{-i}$, $i=0,1,$..., with the postcritical set $P_{n}$
Example 2.3.
Let $f(z)=z^{2}+c$ be infinitely renormalizable. Assume that $f$ is robust and has no unbranched complex bounds ([9], and see below). Then a limit BI map naturally appears as follows. First, at each period $n$ of simple renormalization we have a BI map obtained as in the previous example, where we start with a geodesic $\gamma_{n}$ of the Riemann surface $M$ which is homotopic to a loop in $M$ around $J_{n}$. We rescale this BI map around zero by a linear map $A_{n}$ in such a way that the rescaling postcritical set $P_{n}^{*}=A_{n}(P_{n})$ becomes of size $1$ and obtain, for each such $n$, a rescaled BI map $g_{n}:U_{-i-1}^{(n)}\to U_{-i}^{(n)}$, $i=0,1,2,$..., with the postcritical set $P_{n}^{*}$, where $g_{n}=A_{n}\circ f^{n}\circ A_{n}^{-1}$ is the rescaling of $f^{n}$.
Now, assume that $L_{-}\leq\ell(\gamma_{n})\leq L_{+}$, for some $0<L_{-}<L_{+}<\infty$ and all $n$ along a subsequence $\mathcal{R}$ of simple renormalizations, where $\ell$ measures the hyperbolic length on $M$.
Such maps exists, see [3].
Following technique of [9] (see particularly Theorem 10.12 there), we can show that, for every $i\geq 0$, the sequence of maps $\{g_{n}:U_{-i-1}^{(n)}\to U_{-i}^{(n)}\}_{n\in\mathcal{R}}$ is pre-compact in the Caratheodory topology. It allows us, inductively for each $i=0,1,$..., to pass to a limit along a subsequence of $n$’s and obtain a limit sequence of proper degree two maps $g:U_{-i-1}\to U_{-i}$, $i=0,1,2,$..., alone with a compact $X$ which is a Hausdorff limit point of the sequence $\{P_{n}^{*}\}$. Moreover, $X\subset\cap_{i\geq 0}U_{-i}$, $0\in X$ and, for every $k=1,2,$..., $g_{n}^{k}(0)\to g^{k}(0)\in X$. This implies that all maps $g:U_{-i-1}\to U_{-i}$ share the same postcritical set which is a subset of $X$. Thus $g:U_{-i-1}\to U_{-i}$ ($i\geq 0$) is a BI map.
3. On BI maps
Lemma 3.1.
Let $Y\subset\cap_{i\geq 0}U_{-i}$ be a compact set. The following conditions are equivalent:
  1. (1)
    $g^{-1}(Y)=Y$, i.e., $Y$ is completely invariant,
  2. (2)
    $g_{i}^{-1}(Y)=Y$ for all $i\geq 0$,
  3. (3)
    $g_{i}^{-1}(Y)=Y$ for some $i\geq 0$ and $g_{j}^{-1}(x)=g_{i}^{-1}(x)$ for any $j$ and any $x\in Y$,
  4. (4)
    $g:Y\to Y$ has degree $d$ (counting multiplicities).
Moreover, if either of (1)-(4) holds, then, given $n>0$, for every neighborhood of $W$ of $Y$ there exists another neighborhood $W^{\prime}$ of $Y$ as follows: for every j=1,,nj=1,...,n, two (multi-valued) maps $g_{0}^{-j}=g_{j-1}^{-1}\circ...\circ g_{0}^{-1}$ and $(g_{0}^{-1})^{j}=g_{0}^{-1}\circ...\circ g_{0}^{-1}$ ($j$ times) are well-defined on $W^{\prime}$, $g_{0}^{-j}=(g_{0}^{-1})^{j}$ on $W^{\prime}$ (as multi-valued maps), and $g_{0}^{-j}(W^{\prime})=(g_{0}^{-1})^{j}(W^{\prime})\subset W$.
Proof.
[…]
Let’s verify the last claim. Let $W$ be a neighborhood of $Y$. Fix $j$. By item (3), $g_{0}^{-j}(x_{0})=(g_{0}^{-1})^{j}(x_{0})$ for $x_{0}\in Y$ and all preimages are also in $Y$, hence, this holds for all $x$ in a neighborhood $W_{x_{0}}$ of $x_{0}$ (with all preimages of $x$ in neighborhoods of corresponding preimages of $x_{0}$ and in $W$). By compactness of $Y$, the union of finitely many $W_{x_{0}}$ form a neighborhood $W^{\prime}_{j}$ of $Y$. Then $W^{\prime}=\cap_{j=0}^{n-1}W^{\prime}_{j}$ is as required.
Next, we are going to describe $3$ alternatives for a BI map. Let $g:U^{\prime}\to U$ be a BI map, where $U^{\prime}=\cup_{i=1}^{\infty}U_{-i}$, $U=\cup_{i=0}^{\infty}U_{-i}$. Let us fix a Jordan curve $\Gamma_{0}$ in $U_{0}$ which encloses the postcritical set $P_{g}$. Let $V_{0}$ be the Jordan domain bounded by $\Gamma_{0}$, so that $\overline{V_{0}}\subset U_{0}$.
Recall that $g_{i}^{-n}:U_{-i}\to U_{-i-n}$ is the inverse (multi-values) function to $g^{n}:U_{-i-n}\to U_{-i}$.
Let, for each $n>0$, $V_{-n}=g_{0}^{-n}(V_{0})$ (which is compactly contained in $U_{-n}$) and $\Gamma_{n}=\partial V_{-n}$ (which is also a Jordan curve).
A priori, there are $3$ possibilities:
(A) $\Gamma_{n}\cap\Gamma_{0}\neq\emptyset$ for all $n$.
(B) For some $n>0$, $\overline{V}_{-n}\subset V_{0}$.
(C) For some $n>0$, $\overline{V}_{0}\subset V_{-n}$.
Proposition 3.2.
Assume that (A) holds. Then, for every $\varepsilon>0$ and every $n$ big enough, the map $g^{n}:U_{-n}\to U_{0}$ has a repelling fixed point in $\varepsilon$-neighborhood of $\Gamma_{0}$.
Assume that (B) holds. Then $g^{n}:V_{-n}\to V_{n}$ is a PL map of degree $d^{n}$. Assume additionally that there is a full compact $K\subset\cap_{i\geq 0}U_{-i}$ as in Theorem 1.1, i.e., $K\supset P_{g}$ and $g^{-1}(K)=K$. Then there exist a polynomial-like restriction $g:V^{\prime}\to V$ where $V$ is a simply connected neighborhood of $K$ and $K$ is the filled Julia set of $g:V^{\prime}\to V$.
Assume (C) holds. Then there is an attracting fixed point $a\in\cap_{i\geq 0}g^{-i}(V_{0})$ such that $g^{n}(z)\to a$ for all $z\in\cup_{i\geq 0}g^{-i}(V_{0}$. Moreover, $g$ has no a completely invariant compact. Finally, if the alternative (C) holds for a sequence of $\Gamma_{0}$’s that tends to $\partial U$, then the BI map $g:U^{\prime}\to U$ is trivial: there is an attracting fixed point $a\in\cap_{i\geq 0}U_{-i}$ such that $g^{n}(z)\to a$ for all $z\in U$.
Proof.
(A). In this case, there exist a sequence of indices $n_{i}\to\infty$ and a sequences of points $x_{i}\in\Gamma_{n_{i}}\cap\Gamma_{0}$ as follows: $x_{i}\to b^{\prime}$, $y_{i}:=g^{n_{i}}(x_{i})\in\Gamma_{0}$, $y_{i}\to b$ and $(g^{n_{i}})^{\prime}(x_{i})\to L$ where $L\in\mathbb{C}\cup\{\infty\}$. Fix a small disk $B$ centered at $b$ which is contained in $U_{0}$ and disjoint with $P_{g}$, and, for each $i$, let $\phi_{i}$ be a branch of $g_{0}^{-n_{i}}$ which is defined in $B$ and maps $y_{i}$ to $x_{i}$. Then $(\phi_{i})^{\prime}(y_{i})\to 1/L$.
Thus $\phi_{i}:B\to\mathbb{C}$, $i=1,2,$..., is a sequence of univalent maps and is a normal family because it avoids taking values in the set $P_{g}\cup g^{-1}(P_{g})$, which must contain at least $2$ points by Definition 1.1. By this reason, the case $L=0$ is impossible.
Consider the case $L\neq\infty$, that is, $0<|L|<\infty$. Replacing $(n_{i})$ by a subsequence, in this case $(\phi_{i})$ tends uniformly on compacts in $B$ to a non-constant, hence, univalent function $\nu$ which maps $B$ to a neighborhood $B^{\prime}$ of $b^{\prime}$. As $\psi_{i}(B)\subset g_{0}^{-n_{i}}(U_{0})=U_{-n_{i}}$, then $B^{\prime}\subset U_{-n_{i}}$ for all $i$ big enough. Shrinking slightly $B^{\prime}$, we get:
Claim. The sequence $g^{n_{i}}:B^{\prime}\to B$ is well-defined and converges uniformly to $\nu^{-1}$.
Let $m_{i}=n_{n+1}-n_{n_{i}}$ Passing again to a subsequence of $(n_{i})$, one can assume that $m_{i}\to\infty$. Write $g^{n_{i+1}}=g^{m_{i}}\circ g^{n_{i}}$, on $B^{\prime}$. Here, $g^{n_{i}}(B^{\prime})\subset g^{n_{i}}(U_{-n_{i+1}})\subset U_{-m_{i}}$.
Shrinking again $B^{\prime}$ if necessary, by the Claim, $g^{m_{i}}\to\operatorname{Id}$, uniformly on $\nu^{-1}(B^{\prime})$ where $\nu^{-1}(B^{\prime})\subset U_{-m_{i}}$. Concluding this argument (and shrinking $B$ to a smaller disk around $b$), we obtain that, for some branches $I_{i}:=g_{0}^{-m_{i}}$ on $B$, $I_{i}\to\operatorname{Id}$ uniformly on $B\subset U_{0}$.
Let now $D$ be any simply connected subdomain of $U_{0}\setminus P_{g}$ containing $B$. As all branches of $g_{0}^{-m_{i}}$ are well-defined in $D$ and form a normal family, passing to a subsequence, we get that $I_{i}$ (more precisely, their extensions from $B$ to $D$) converge uniformly on compacts in $D$ to $\operatorname{Id}$. As each $I_{i}$ is an inverse to $g^{m_{i}}$, it follows that, given a compactly contained in $D$ domain $F\subset D$, for all $i$ big enough, we have: $F\subset U_{-m_{i}}$ and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $F$.
We apply this to two domains $D_{1},D_{2}$ which together cover the curve $\Gamma_{0}$, and to their compactly contained subdomains $F_{1}\subset D_{1}$, $F_{2}\subset D_{2}$, such that $F_{1}\cup F_{2}\supset\Gamma_{0}$. As a result, we obtain that $\Gamma_{0}\subset U_{-m_{i}}$ for all $i$ big enough and $g^{m_{i}}\to\operatorname{Id}$ uniformly on $\Gamma_{0}$. This is a contradiction by the following reason: on the one hand, as we have just shown, $\Gamma_{0}\subset U_{-m_{i}}$, also $\Gamma_{0}=\partial V_{0}$ and $g^{m_{i}}(U_{-m_{i}})=U_{0}$, hence, the sequence $g^{m_{i}}$ is well-defined on $\overline{V_{0}}$ and converges uniformly to the identity map $\operatorname{Id}$. Then $(g^{m_{i}})^{\prime}\to 1$ uniformly on compacts in $V_{0}$. On the other hand, $(g^{m_{i}})^{\prime}(c)=0$ for each $c\in C_{g}$ This contradiction excludes the case $0<|L|<\infty$.
Thus we must have that $(g^{n_{i}})^{\prime}(x_{i})\to\infty$, i.e., $(\psi_{i})^{\prime}(y_{i})\to 0$. Let us show that then $g^{n_{i}}$ must have a repelling fixed point close to the curve $\Gamma_{0}$. Indeed, let $D\subset U_{0}\setminus P_{g}$ be a simply connected domain that contains points $b^{\prime}=\lim x_{i}$ and $b=\lim y_{i}$. Then, by the normality of the family of extensions of $\phi_{i}$ from $B$ to $D$, a subsequence of $\phi_{i}$ converges uniformly on compacts in $D$ to a point $A\in D$. In particular, for a small disk $B_{A}$ centered at $A$, and for some branch $\phi_{i}$ of $g^{-n_{i}}$, $\overline{\phi_{i}(B_{A})}\subset B_{A}$. Therefore, $\phi_{i}$ has an attracting fixed point $a\in B_{A}$, which is thus a repelling fixed point of $g^{n_{i}}$.
(B). $\overline{V}_{-n}\subset V_{0}$ for some fixed $n>0$. Then $G:=g^{n}:V_{-n}\to V_{0}$ is a polynomial-like map (of degree $d^{n}$). Let $K_{G}$ be the filled Julia set of $G$. Since $g^{-1}(K)=K$ and using Lemma 3.1, $G^{-1}(K)=g_{0}^{-n}(K)=K$. As $G:V_{-n}\to V_{0}$ is conjugate to a polynomial and $K\geq 2$, it follows that $K=K_{G}$. Moreover, as $P_{G}=P_{g}\subset V_{0}$, hence, $K=K_{G}$ is connected.
It remains to show $g$ itself has a PL restriction, i.e., there exist a simply connected neighborhood $V$ of $K$, such that $V^{\prime}:=g^{-1}(V)$ is compactly contained in $V$. If $n=1$, there is nothing to prove (set $V=V_{0}$).
So let $n\geq 2$. We present two proofs: the first one uses heavily that $K$ is a continuum while the second one is "continuum free" and can be easily adapted to the set up of Theorem 1.3 where $K$ is disconnected.
First proof. The following argument is well-known. Let $\Psi:\mathbb{C}\setminus K\to\mathbb{D}^{*}:=\{|w|>1\}$ be the conformal isomorphism such that $\Psi(\infty)=\infty$. Below, $\hat{A}=\Psi(A\setminus K)$ for a neighborhood $A$ of $K$, and $A_{*}$ is the union of $\hat{A}$ with its symmetric w.r.t. $S^{1}$ completed by $S^{1}$. Similar notations are used for maps. Replace $g:U_{-1}\setminus K\to U_{0}\setminus K$ by $\hat{g}:=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-1}\to\hat{U}_{0}$. By Schwartz’s reflexion principle, $\hat{g}$ extends to a holomorphic map $g_{*}:(U_{-1})_{*}\to(U_{0})_{*}$. The latter map has no critical points as $g_{*}^{-1}(S^{1})=S^{1}$. Let $\hat{G}=\Psi\circ g\circ\Psi^{-1}:\hat{U}_{-m}\to\hat{U}_{0}$ and $G_{*}:(U_{-m})_{*}\to(U_{0})_{*}$. As this map is a covering and $\overline{(U_{-m})_{*}}\subset(U_{0})_{*}$, then $G_{*}$ expands the hyperbolic metric $\rho$ of the domain $(U_{0})_{*}$, i.e, $D_{G}(w):=\rho\circ G(w)|G^{\prime}(w)|/\rho(w)>1$ for $w\in(U_{-m})_{*}$. Now note that $G_{*}=g_{*}^{m}$ in some neighborhood of $S^{1}$. Define a new metric $\tilde{\rho}=\sum_{i=0}^{m-1}\rho\circ g_{*}^{i}|(g_{*}^{i})^{\prime}|$. Then $\tilde{D}_{g_{*}}(w):=\frac{\tilde{\rho}\circ g_{*}(w)|g_{*}^{\prime}(w)|}{\tilde{\rho}(w)}=1+\frac{\rho(w)}{\tilde{\rho}(w)}\{D_{G}(w)-1\}>1$ for all $w$ near $S^{1}$. By the compactness, $\tilde{D}_{g_{*}}(w)\geq\lambda$ for some $\lambda>1$ and all $w$ in a small annulus around $S^{1}$. It follows that there is a neighborhood $V_{*}$ of $S^{1}$ which is symmetric w.r.t. $S^{1}$ such that $g_{*}^{-1}(V_{*})$ is compactly contained in $V_{*}$. Then we let $V=\Psi^{-1}(V_{*}\cap\mathbb{D}^{*})\cup K$.
Second proof. Here we don’t use that $K$ is a continuum.
As $V_{-n}=g_{0}^{-n}(V_{0})$ and $\overline{V_{-n}}\subset V_{0}$, then the iterates $(g_{0}^{-n})^{j}$ ($j>0$) of $g_{0}^{-n}$ are all well-defined on $V_{0}$ and $\tilde{V}_{j}:=(g_{0}^{-n})^{j}(V_{0})$, $j=0,1,$..., form a decreasing sequence of Jordan domains (with $\tilde{V}_{0}=V_{0}$)such that
(3.1) $K=K_{G}=\cap_{j>0}(\tilde{V}_{j}).$
Note that $\tilde{V}_{j+1}=g_{0}^{-n}(\tilde{V}_{j})$ and $\overline{\tilde{V}_{j+1}}\subset\tilde{V}_{j}$.
For every $j\geq 0$, let $\Omega_{j}:=\cup_{i=0}^{n-1}g_{0}^{-i}(\tilde{V}_{j})$.
Apply the last claim of Lemma 3.1 to a neighborhood $W=V_{0}$ of $K$ and find a neighborhood $W^{\prime}$ of $K$ such that
(3.2) $g_{0}^{-i}=(g_{0}^{-1})^{i}\mbox{ for }i=0,1,...,n.$
By (3.1), there exists $j_{*}>0$ such that, for every $j\geq j_{*}$, $K\subset\Omega_{j}\subset W^{\prime}$ and $\Omega_{j}$ shrink to $K$ as $j\to\infty$.
This allows us to write, using (3.2), for every $j\geq j_{*}$:
$g_{0}^{-1}(\Omega_{j})=g_{0}^{-1}(\cup_{i=0}^{n-1}(g_{0}^{-1})^{i}(\tilde{V}_{j}))=\cup_{i=1}^{n}(g_{0}^{-1})^{i}(\tilde{V}_{j})=$
$\Omega_{j}\cup[(g_{0}^{-1})^{n}(\tilde{V}_{j})\setminus\tilde{V}_{j})]=\Omega_{j}\cup[(g_{0})^{-n}(\tilde{V}_{j})\setminus\tilde{V}_{j})]=$
$\Omega_{j}\cup\tilde{V}_{j+1}\setminus\tilde{V}_{j}\subset\Omega_{j}.$
Thus $g_{0}^{-1}(\Omega_{j})\subset\Omega_{j}$ and
$g:g_{0}^{-1}(\Omega_{j})\setminus K\to\Omega_{j}\setminus K$
is an unbranched covering map, for each $j\geq j_{*}$.
Let $\Omega^{*}=\Omega_{j_{*}}\setminus K$ and let $\rho_{*}$ be the hyperbolic metric of the domain $\Omega^{*}$. As $g_{0}^{-1}(\Omega^{*})\subset\Omega^{*}$ and there is no an equality, by the above and the Schwartz lemma, every local branch of $g_{0}^{-1}:\Omega^{*}\to\Omega^{*}$ is a strict contraction.
Now, fix $j$ large enough, such that $\overline{\Omega_{j}}\subset\Omega_{j_{*}}$.
Let $Y$ be the topological hull of $\overline{\Omega_{j}}$. As $g_{0}^{-1}(\overline{\Omega_{j}})\subset\overline{\Omega_{j}}$ and $g:U_{-1}\to U_{0}$ is a proper map between simply connected domains, $g_{0}^{-1}(Y)\subset Y$. Hence, if $Y^{*}=Y\setminus K$, then $g_{0}^{-1}(Y^{*})\subset Y^{*}$ as well (remind that $K$ is completely invariant).
The set $\Delta:=Y\setminus Int(g_{0}^{-1}(Y))$ is a compact subset of $\Omega^{*}$, therefore, there exist $r>0$ and $\kappa\in(0,1)$ as follows: for every $a\in\Delta$, $B_{h}(a,2r)$ is a topological disk and every branch $g_{0}^{-1}:B_{h}(a,2r)\to\Omega_{*}$ contracts the distance in $\Omega^{*}$ by a factor at leat $\kappa$ (here and below in the proof $B_{h}(x,t)$ is the disk centered at $x\in\Omega^{*}$ and of radius $t>0$ in the hyperbolic metric $\rho_{*}$ of $\Omega^{*}$).
Note that $\partial Y\subset\partial\Omega_{j}\subset\Omega^{*}$ and $\partial Y^{*}$ is the disjoint union of $\partial Y$ and $\partial K$. Let $Y^{*}_{r}=Y^{*}\cup_{a\in\partial Y}\bar{B_{h}}(a,r)$.
Let $\Gamma_{*}$ be the outer boundary of $Y^{*}_{r}$ (i.e, the boundary of the topological hull of $Y^{*}_{r}$). Note that $\Gamma_{*}$ is a compact subset of $\Omega^{*}$. Let $V_{*}$ be a (simply connected) domain bounded by $\Gamma_{*}$. We claim that $g_{0}^{-1}(V_{*})$ is compactly contained in $V_{*}$. It’s enough to show that $g_{0}^{-1}(\Gamma_{*})\subset V_{*}$. To this end, let $z\in\Gamma_{*}$ and $z^{\prime}\in g_{0}^{-1}(z)$. There exist $a\in\partial Y$, such that $\rho_{*}(z,a)\leq r$, and a branch $\psi$ of $g_{0}^{-1}$ in a neighborhood of $\bar{B}_{h}(a,r)$ such that $\psi(z)=z^{\prime}$. If $a^{\prime}:=\psi(a)$, then $\rho_{*}(a^{\prime},z^{\prime})\leq k\rho_{*}(a,z)\leq kr<r$. As $a^{\prime}\in g_{0}^{-1}(\partial Y)\subset\Delta$, it follows that $z^{\prime}\in Y^{*}_{r}$. Thus $V^{\prime}_{*}:=g_{0}^{-1}(V_{*})$ is compactly in $V_{*}$, i.e., $g:V_{*}^{\prime}\to V*$ is the required PL map.
(C). If $\overline{V}_{0}\subset V_{-n}$ for some $n>0$, then $\overline{g^{n}(V_{0})}\subset V_{0}$. Hence, by the Wolff-Denjoy theorem (e.g., [2]), the sequence of domains $g^{kn}(V_{0})$, $k=1,2,$..., shrink to an attracting fixed point $a\in V_{0}$ of $g^{n}$. Then $g^{i}(\overline{V_{0}})$, $i\geq n$, tend to the cycle $O\ni a$. As $P_{g}\subset V_{0}$ is a closed invariant set, then $O\subset P_{g}\subset V_{0}$. Therefore, $O=g^{i}(O)\subset g^{i}(V_{0})$ for all $i\geq n$ where $g^{nk+j}(V_{0})$ shrinks to $g^{j}(a)$ as $k\to\infty$. Since $P_{g}\subset g^{i}(V_{)})$ for all $i\geq n$, it follows that $O=\{a\}$. If a completely invariant compact $K\subset V$ exists, then $K=\{a\}$ and $f^{-1}(a)=a$. In other words $P_{g}=\{a\}$ and $a$ is a fixed critical point of order $d$, which is excluded by Definition 1.1. Finally, if the alternative (C) holds for a sequence of such $\Gamma_{0}$’s that tends to $\partial U_{0}$, then the sequence of corresponding Jordan domains $V_{0}$’s bounded by $\Gamma_{0}$’s exhausts $U_{0}$, hence, by the above, there is an attracting fixed point $a\in\cap_{i\geq 0}g^{-i}(U_{0})$ such that $g^{n}(z)\to a$ for all $z\in U=\cup_{i\geq i}U_{0}$, i.e., $g:U^{\prime}\to U$ is trivial.
4. Proof of Theorem 1.2
Assume the contrary: $g:U^{\prime}\to U$ has a completely invariant full compact set $K$ such that $P_{g}\subset K\subset\cap_{n\geq 0}U_{-n}$, and there exist a point $a\in U_{0}\setminus K$ and sequences $n_{i}\to\infty$ and $a_{i}\to a$ such that $a_{i}\in U_{-n_{i}}$, $g^{n_{i}}(a_{i})=a_{i}$ and, for each $i$, $|\lambda_{i}|>1$ where $\lambda_{i}=(g^{n_{i}})^{\prime}(a_{i})$.
For every $i$, as $a_{i}\in U_{0}\cap U_{-n_{i}}$ and $|\lambda_{i}|>1$, there exists a branch $\psi_{n_{i}}$ of $g_{0}^{-n_{i}}$ which is defined (hence, univalent) in some neighborhood of $a_{i}$ such that $\psi(a_{i})=a_{i}$. As $a\notin P_{g}$, there exists a disk $Z:=B(a,r)\subset U_{0}$ such that $\overline{Z}\subset U_{0}\setminus P_{g}$. Therefore, each $\psi_{i}$ extends to a univalent map to $Z$. Repeating arguments as in the proof of the alternative (A), Proposition 3.2 (where one can take $V_{0}$ to be any Jordan domain such that $P_{g}\subset V_{0}\subset\overline{V_{0}}\subset U_{0}$ and $a\in\Gamma_{0}:=\partial V_{0}$) we obtain that $\lambda_{i}\to\infty$. Equivalently, $\psi_{i}^{\prime}(a_{i})=\lambda_{i}^{-1}\to 0$. In turn, shrinking the radius $r$ of the disk $Z$ if necessary, this implies that $a_{i}\in\psi_{n_{i}}(Z)\subset\overline{\psi_{n_{i}}(Z)}\subset Z$ for all $i$ big enough.
4.1. The case of connected $K$
Let $\pi:\mathbb{D}^{*}:=\{|w|>1\}\to\mathbb{C}\setminus K$ be the conformal isomorphism such that $\pi(\infty)=\infty$. Replace $g:U^{\prime}\setminus K\to U\setminus K$ by $\hat{g}:=\pi^{-1}\circ g\circ\pi:\hat{U}^{\prime}\to\hat{U}$ where $\hat{U}^{\prime}=\pi^{-1}(U^{\prime}\setminus K)$, $\hat{U}=\pi^{-1}(U\setminus K)$. By Schwartz’s reflection principle, $\hat{g}$ extends to a holomorphic map $\hat{g}_{*}:\hat{U}^{\prime}_{*}\to\hat{U}_{*}$ where $E_{*}$ denotes the union of a domain $E\subset\mathbb{D}^{*}$ containing $S^{1}$ in its boundary with its symmetric w.r.t. $S^{1}$ completed by $S^{1}$.
On the other hand, $\hat{g}_{*}^{-1}(S^{1})=S^{1}$, hence $\hat{g}_{*}$ has no critical points. Therefore, all pullbacks of $\hat{g}_{*}^{n}$ are well-defined in any simply connected subdomain of $(\hat{U}_{0})_{*}$. Fix such a subdomain $D$ that contains $\hat{a}$ and an arc $l$ of $S^{1}$. As $\hat{a}_{i}\to\hat{a}$ and, for each $i$, $\hat{a}_{i}$ is an attracting fixed point of the map $\hat{\psi}_{i}:\hat{Z}\to\mathbb{C}$ and $\hat{\psi}_{i}^{\prime}(\hat{a}_{i})\to 0$, we arrive at a sequence of holomorphic extensions $(\hat{\psi}_{i})_{*}$ of $\hat{\psi}_{i}$ to $D$ such that $(\hat{\psi}_{i})_{*}\to\hat{a}$ in $D$. But this cannot hold for points of the arc $l$, by the complete invariance of $S^{1}$, a contradiction
The proof for the disconnected $K$ requires new ideas and is substantially more technical. It also holds in the case of connected $K$. Let’s indicate an idea returning to the connected $K$. We don’t consider the reflection of the map $\hat{g}:\hat{U}^{\prime}\to\hat{U}$ across $S^{1}$. Instead, let us fix a simply connected domain $\Omega$ in $\mathbb{D}^{*}=\{|z|>1\}$ such that, if $\hat{Z}=\pi^{-1}(Z)$, then $\hat{a}\in\hat{Z}\subset\Omega$ and $\partial\Omega\cap S^{1}$ contains a nontrivial arc $l$ of $S^{1}$ (in fact, the set $\partial\Omega\cap S^{1}$ having a positive length would be enough). Since all branches $\hat{g}_{0}^{-n}$ are well-defined in $\Omega$, for every $i$, the local map $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ extends to a univalent map $\hat{\psi}_{i}:\Omega\to\mathbb{C}$ where $\hat{\psi}_{i}(\hat{a}_{i})=\hat{a}_{i}$ and $\hat{\psi}_{i}^{\prime}(\hat{a}_{i})\to 0$. The point is that this implies that $\hat{\psi}_{i}$ expands the length of the arc $l$ by factors tending to $\infty$, see Appendix C. This is clearly a contradiction since all images of $l$ must lie in $S^{1}$.
4.2. Proof of Theorem 1.1 for disconnected $K$
By Theorem A.1 and Corollary A.1, the logarithmic capacity of the completely invariant compact set $K$ is positive so Corollary B.2 applies.
Let us fix an open subset of $\mathbb{C}$ as follows:
- it contains the fixed point $a$,
- it consists of finitely many, but at least two, Jordan domains $Q_{1},...,Q_{m}$, $m\geq 2$, such that $\overline{Q}_{i}$, i=1,,mi=1,...,m, are pairwise disjoint and $K\subset\cup_{i=1}^{m}Q_{i}$.
Let $\varepsilon>0$ be the distance between the set $\cup_{i=1}^{m}\partial Q_{i}$ and $K$.
For this $\varepsilon$, we find a simply connected domain $W$ as in Corollary B.2, in particular, the connected set $\pi(W)\subset\mathbb{C}\setminus K$ must lie in one and only one of the domains $Q_{1},...,Q_{m}$. W.l.o.g. one can assume that
$\overline{\pi(W)}\subset Q_{1}.$
Let us joint the point $a$ and a point $q\in\partial\pi(W)\setminus K$ by a simple arc $\alpha$ which, on the one hand, lies in the range $U_{0}$ of $g$, and, on the other hand, except its end point $q$, lies outside of $\overline{\pi(W)\cup K}$ as well as outside of all others $\overline{Q_{i}}$, $2\leq i\leq m$, and also crosses the Jordan curve $\partial Q_{1}$ at a unique point. Choose a point $\hat{q}\in\partial W\cap\mathbb{D}$ such that $\pi(\hat{q})=q$ and let $\hat{\alpha}$ be the lift of the curve $\alpha$ by the covering $\pi$ that begins at $\hat{q}$. Then $\tilde{\alpha}$ is also a simple arc (in $\mathbb{D}$) with the end points $\hat{q}$ and some $\hat{a}\in\mathbb{D}$ such that $\pi(\hat{a})=a$. As $\pi$ is a local homeomorphism near $\hat{a}$, one can choose $r>0$ small enough so that there exists a topological disk $\hat{Z}$ around $\hat{a}$ such that $\pi:\hat{Z}\to Z$ where $Z=B(a,r)$ is a homeomorphism.
Note that $\hat{\alpha}\cap\overline{W}=\{\hat{q}\}$. Let us slightly "thicken" the curve $\hat{\alpha}$ to a narrow "tube" $T$ so that $\Omega:=\hat{Z}\cup T\cup W$ is a simply connected domain in $\mathbb{D}$. Then $\pi(\Omega)=Z\cup\pi(T)\cup\pi(W)$ is a domain in $\mathbb{C}\setminus K$ which is disjoint with $\overline{Q_{i}}$, for all $2\leq i\leq m$. As $\alpha$ crosses $\partial Q_{1}$ at a single point, we can also assume that $Q:=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with all others $Q_{2},...,Q_{m}$.
For each $i$ big enough, we can "lift" the map $\psi_{i}:Z\to Z$ to a univalent map $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ so that $\pi\circ\hat{\psi}_{i}=\psi_{i}\circ\pi$ on $\hat{Z}$.
Claim 0. $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ extends to a well-defined holomorphic function (also called $\hat{\psi}_{i}$) in the domain $\Omega$.
Indeed, let $\hat{\beta}$ be any curve in $\Omega$ that starts at $\hat{a}$. Then $\beta:=\pi\circ\hat{\beta}$ is a curve in $\pi(\Omega)$ that starts at $a$ and lies in $U_{0}\setminus K$. As $g^{n_{i}}:U_{-n_{i}}\setminus K\to U_{0}\setminus K$ is an unbranched covering map, the local branch $\psi_{i}:Z\to Z$ of $g^{-n_{i}}$ has an analytic continuation along the curve $\beta$, and it is lifted by the unbranched covering map $\pi:\mathbb{D}\to\hat{\mathbb{C}}\setminus K$ to an analytic continuation of $\hat{\psi}_{i}$ along the curve $\hat{\beta}$. Since the latter is any curve starting at $\hat{a}$ in the simply connected domain $\Omega$, by the Monodromy theorem, the function $\hat{\psi}_{i}$ is well-defined holomorphic continuation of $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ to the whole domain $\Omega$.
Moreover,
Claim 1. For each $i$, the function $\hat{\psi}_{i}:\Omega\to\mathbb{C}$ is univalent.
Indeed, assume the contrary: $\hat{\psi}_{i}(x_{1})=\hat{\psi}_{i}(x_{2})$ for some $x_{1}\neq x_{2}$ in $\Omega$. Let $\hat{l}$ be a simple curve in $\Omega$ joining $x_{1}$ and $x_{2}$. As all branches of $g^{-n_{i}}$ are inverse branches of a well-defined map $g^{n_{i}}:U_{-n}\to U_{0}$ and since $Q=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with some other $Q_{j}$, $j\neq 1$, the following must hold:
  • $\pi(x_{1})=\pi(x_{2})$, i.e., $l:=\pi\circ\hat{l}$ is a closed curve in $\pi(\Omega)\subset Q\setminus K$,
  • an analytic continuation of some local branch of $g^{-n_{i}}$ along the closed curve $l$ maps this curve again onto some closed curve $L$,
  • there exists a lift of $L$ by $\pi$ to a curve $\hat{L}$ which begins and ends at the same point $\hat{y}:=\hat{\psi}_{n}(x_{1})=\hat{\psi}_{n}(x_{2})$, in other words the lift of $L$ is again a closed curve.
The latter condition implies that $L$ must be homotopic to a point in the domain $\hat{\mathbb{C}}\setminus K$. On the other hand, as the curve $l$ lies in a simply connected domain $Q$ which is disjoint with some $Q_{j}$, the curve $L$ must lie in one of the components of $g^{-n_{i}}(Q)$, which, in turn, are disjoint with all components of $g^{-n_{i}}(Q_{j})$. As each of the latter components contains points of $K$, the curve $L$ cannot enclose $\infty$ (in $\hat{\mathbb{C}}$), hence, $L$ is homotopic to a point in the subdomain $\mathbb{C}\setminus K$ of $\mathbb{C}$. Since $K$ is completely invariant, then $l$ is homotopic to a point in $\mathbb{C}\setminus K$ too, hence, its lift $\hat{l}$ must be a closed curve, a contradiction. This proves Claim 1.
Let $F=\partial W\cap\partial\mathbb{D}$. Thus, we have:
  1. ($\hat{1}$)
    for every big $i$, $\hat{\psi}_{i}:\Omega\to\mathbb{D}$ is a univalent continuation of $\hat{\psi}_{i}:\hat{Z}\to\hat{\mathbb{Z}}$,
  2. ($\hat{2}$)
    $\hat{\psi}_{i}(\hat{a}_{i})=\hat{a}_{i}$, $\hat{a}_{i}\to\hat{a}\in\Omega$ and $\hat{\psi}_{i}^{\prime}(\hat{a}_{i})=\psi_{i}^{\prime}(a_{i})\to 0$,
  3. ($\hat{3}$)
    as $|F|>0$ and $K$ is completely invariant, for every big $i$ and almost every $w_{0}\in F$, $\hat{\psi}_{i}(w)$ has a an angular limit $\hat{\psi}(w_{0})$ as $w\to w_{0}$ in $\Omega$ and $\hat{\psi}_{i}(w_{0})\in\partial\mathbb{D}$.
We claim that ($\hat{1}$)-($\hat{3}$) are incompatible.
Indeed, let $R:\mathbb{D}\to\Omega$ be a Riemann map such that $R(0)=\hat{a}$. Let $f_{i}=\hat{\psi}_{i}\circ R$, a univalent map in $\mathbb{D}$, for each $i$. As $\partial\Omega$ is a rectifiable Jordan curve, $R^{\prime}\in H^{1}(\mathbb{D})$. Hence, $E:=R^{-1}(F)$ is a measurable subset of $S^{1}$ of a positive 1-dim measure $|E|>0$.
Let $f_{i}=\hat{\psi}_{i}\circ R$, a univalent map of $\mathbb{D}$ onto a bounded domain in $\mathbb{D}$, for each $i$. Since $\hat{\psi}_{i}:\Omega\to\mathbb{D}$ are univalent and form a normal family, by ($\hat{2}$), $f_{i}(0)=\hat{\psi}_{i}(\hat{a})\to\hat{a}$ where $\hat{a}\in\Omega\subset\mathbb{D}$, and $f_{i}^{\prime}(0)=\hat{\psi}_{i}^{\prime}(\hat{a})R^{\prime}(0)\to 0$. By ($\hat{3}$), $f_{i}(E)\subset S^{1}$, for all $i$. Therefore, one can choose $i_{*}$ such that
$|E|\sqrt{\frac{(1-|f_{i_{*}}(0)|)^{3}}{(1+|f_{i_{*}}(0)|)|f_{i_{*}}^{\prime}(0)|}}>2\pi,$
in a contradiction with Corollary C.3 applied to $f=f_{i_{*}}$.
5. Proof of Theorem 1.1
It now follows directly from Theorem 1.2 along with Proposition 3.2, where the alternative (A) must occur.
6. Proof of Corollary 1.5 and its Complement
Proof of (b). By Proposition 3.2, for every Jordan curve $\Gamma$ in $U_{0}$ which encloses $P_{g}$ either (A) or (C) must hold. If (C) holds for a sequence of such $\Gamma_{0}$ that tends to $\partial U$, then, by (C), $g:U^{\prime}\to U$ is trivial, which is ruled out by a condition. Therefore, for each Jordan curve $\Gamma$ which surrounds $P_{g}$ and is close enough to $\partial U_{0}$ the alternative (A) holds, that is, $\Gamma_{0}$ must intersect $R^{\prime}$. Now, consider a connected component $C$ of $R_{g}\cup\partial U_{0}$ which contains $\partial U_{0}$ It is enough to prove that it also contains points of $U_{0}$. Assume the contrary, i.e., $C=\partial U$. Fix some Jordan curve $\Gamma^{\prime}$ as above (i.e., $\Gamma^{\prime}$ surrounds $P_{g}$ and close to $\partial U_{0}$) and let $C^{\prime}$ be a connected component of $R_{g}\cup\Gamma^{\prime}$ that contains $\Gamma^{\prime}$. Let $\Omega$ be the unbounded component of $\mathbb{C}\setminus C^{\prime}$ and $D=(\Omega\cap U_{0})\setminus R^{\prime}$. By the assumption, $D$ is an open connected set and $\partial U_{0}$ is a component of the boundary $\partial D$ while $\partial\Omega$ is another component of $\partial D$. Therefore, there exists a curve $\Gamma^{\prime\prime}$ in $D$ which surrounds $\partial\Omega$, hence, $P_{g}$ as well, and $\Gamma^{\prime\prime}$ is close to $\partial U_{0}$. Then $\Gamma^{\prime\prime}$ must intersect $R^{\prime}$, a contradiction because $\Gamma^{\prime\prime}\subset D$ and $D$ is disjoint with $R^{\prime}$.
Proof of (a). Item (1): by Theorem 1.1, $K$ is the filled Julia set of a PL restriction $g:g_{0}^{-1}(W)\to W$. By the Straightening theorem and the Fatou-Julia theory, $\partial K$ is equal to the limit set of repelling periodic points of $g:K\to K$. On the other hand, by Lemma 3.1 applied to $K$, $R_{n}\cap K$ coincides with the set $\{z\in K:g^{n}(z)=z,|(g^{n})^{\prime}(z)|>1\}$. Hence, $R^{\prime}\cap K=\partial K$, i.e., item (1). Item (2): assume, by a contradiction, that $x\in W\setminus K$ and $g^{n}(x)=x$ for some $n\in\mathbb{N}$ and $x\in U_{-n}\cap U_{0}$. On the one hand, $x\in g_{0}^{-n}(x)$. On the other hand, $\overline{g_{0}^{-1}(W)}\subset W$. Let $A_{k}=g_{0}^{-k}(W)\setminus\overline{g_{0}^{-k-1}(W)}$, $k=0,1,2,$.... Then $\{A_{n}\}_{n\geq 0}$ is a partition of $W\setminus K$. If now $x\in A_{j}$, for some $j\geq 0$, then $x\in g_{0}^{-n}(x)\in A_{j+n}$ while $A_{j}\cap A_{j+n}=\emptyset$, a contradiction. Item (3) follows directly from Theorem 1.2.
Proof of the Complement. After the proven Corollary and Remark 1.6, we only need to show the claim about iterates of $x$. Let $x\in V^{\prime}\setminus K$ be such that $O(x)$ is well-defined but $\omega(x)\nsubseteq\partial V$. In other words, there exist a sequence $n_{i}\to\infty$ a compact $E\subset V$ so that $g^{n_{i}}(x)\in E$, for all $i$. Choose a Jordan domain $V_{0}$ such that $P_{g}\cup E\subset V_{0}\subset\overline{V_{0}}\subset V$. As $P_{g}\subset K$, by Theorem 1.2 along with Proposition 3.2, the alternative (B) of Proposition 3.2 must hold, i.e., for some $n\in\mathbb{N}$, $V_{-n}:=g^{-n}(V_{0})$ is contained in $V_{0}$ with its closure. By (B), then $K=K_{G}:=\cap_{k\geq 0}g^{-kn}(V)$. On the other hand, write $n_{i}=l_{i}n+r_{i}$, with integers $l_{i},r_{i}$ such that $l_{i}\to\infty$ and $0\leq r_{i}<n$. Passing to a subsequence, one can assume that $r_{i}=r$ for all $i$. Then $g^{l_{i}n}(g^{r}(x))\in E\subset V_{0}$, for all $i$, hence, $g^{r}(x)\in\cap_{i\geq 0}g^{-l_{i}n}(V_{0})\subset\cap_{k\geq 0}g^{-kn}(V_{0})$, where the latter equality follows from the fact that the sequence of domains $\{g^{-kn}(V_{0})\}_{k\geq 0}$ is decreasing. Thus $g^{r}(x)\in K$ so that $x\in g^{-r}(K)=K$, a contradiction.
7. Proof of Theorem 1.3
The proof follows almost literally the one of Theorem 1.1 for disconnected $K$, where very minor changed are required and only in Proposition 3.2 and its proof. Curves $\Gamma_{n}$ turn into multi-curves. In the alternative (C), $\overline{V_{0}}$ should be a subset of a component of $V_{-n}$. In the part of the proof of Proposition 3.2 where (A) holds, the normality argument in the case $P_{g}=\emptyset$ can be easily amended; in the part where (B) holds, $V$ should be a neighborhood of $K$ consisting of a finitely many simply connected components and the Second proof of the construction of $V$ such that $g^{-1}(V)$ is compactly in $V$ should be chosen.
Appendix A Entropy of locally analytic maps
The following result is contained essentially in [8], [13]:
Theorem A.1.
(1) Let $X\subset\mathbb{C}$ be a compact subset of the plane and $f:X\to X$ be a continuous map which extends to a holomorphic map $f:W^{\prime}\to W$ of some neighborhoods $W^{\prime},W$ of $X$, such that $f^{-1}(X)=X$ and $\deg f|_{X}=N\geq 2$. Assume that $X$ is not equal to the union of cycles of critical points of $f$. Then $h_{top}(f)\geq\log N$, for the topological entropy $h_{top}(f)$ of $f:X\to X$.
(2) If, additionally, $f:W^{\prime}\to W$ is a proper map of degree $N$ where $W$ is simply connected, then the conclusion holds unless $W^{\prime}$ is (simply) connected and $X$ is a singleton $\{c\}$ where $c$ is a fixed critical point of $f$ of multiplicity $N$.
Proof.
(1) We follow closely [13] (where $X$ is the Julia set of a rational function $f$) and [8] (where $f$ is a smooth self-map of a compact smooth manifold) taking into account that our map $f$ is defined only in a small neighborhood of $X$. Shrinking $W$ if necessary one can assume that all critical points of $f:W^{\prime}\to W$ lie in $X$. Let $C$, $C_{s}$ be the sets of non periodic and periodic critical points of $f:X\to X$, respectively. (It is possible though that either $C_{s}$ or both sets $C,C_{s}$ are empty.) $C(\sigma)=\cup_{c\in C}B(c,\sigma)$ be the $\sigma$-neighborhood of $C$ and $C_{s}(\sigma)$ be a neighborhood of $C_{s}$ which is forward invariant ($f(C_{s}(\sigma))\subset C_{s}(\sigma)$) and is contained in the $\sigma$-neighborhood of $C_{s}$ (all distances are Euclidean). Note that $C(\sigma)$ is disjoint with $C_{s}(\sigma)$ for all $\sigma>0$ small enough. We prove:
($\star$) given $t\in(0,1)$, there exist $\sigma>0$ small enough and $k_{0}$ such that for every $k\geq k_{0}$ and every $y\in K$, if $y,f(y),f^{2}(y),...,f^{k}(y)\notin C_{s}(\sigma)$ then
(A.1) $\{j:0\leq j<k,f^{j}(y)\in C(\sigma)\}<tk.$
First, if $y$ is attracted by an attracting (not superattracting) or parabolic cycle, such that some $c\in C$ is also in the basin of attraction of the same cycle (the number of such cycles is at most #C\#C), then (A.1) holds because the left-hand side of (A.1) is bounded, uniformly over $y$ provided $\sigma$ is small enough. Hence, one can assume that
($\star\star$) $y$ is not attracted by a cycle that attracts also a point of $C$.
In this case, since $C,C_{s}$ are finite sets, (A.1) would follow if we show that
(A.2) $\inf\{m:z,f^{m}(z)\in B(c,\sigma)\mbox{ for some }c\in C,z\in X\}\to\infty$
as $\sigma\to 0$. To this end, let $r>0$ be such that $X_{r}:=\cup_{a\in X}B(a,r)\subset W^{\prime}$ and $M=\sup\{|f^{\prime}(z)|:z\in X_{r}\}<\infty$. Given $c\in C$, there exist $\rho>0$, $L>0$, $l\geq 2$ (the order of $c$) such that $L^{-1}|z-c|^{l-1}<|f^{\prime}(z)|<L|z-c|^{l-1}$ for all $z\in B(c,\rho)$. We prove that there exists $L_{*}$ that depends on $M$, $L$ and $l$ such that $m>L_{*}/\log(1/\sigma)$ whenever $\sigma<\min\{r/8,\rho\}$ and $z,f^{m}(x)\in B(c,\sigma)$ for some $x\in X$. It is enough to prove that
$L(5\sigma)^{l}M^{m}\geq 2\sigma.$
If we assume the contrary, then, on the one hand, all images $f^{j}(B(x,4\sigma))$, j=0,1,2,,mj=0,1,2,...,m, are well-defined, i.e. are contained in $X_{r}\subset W^{\prime}$. On the other hand, as $|x-f^{m}(x)|<2\sigma$, we obtain that $f^{m}(B(x,4\sigma))\subset B(x,4\sigma)$ which implies that $f^{mk}$ tends as $k\to\infty$ to an attracting fixed point of $f^{m}$ uniformly on compacts in $B(x,4\sigma)$. Note that $c\in B(x,4\sigma)$, therefore, $x$ and $c$ are attracted by an attracting cycle of $f$, in contradiction with ($\star\star$). This proves ($\star$).
We proceed by fixing any $\alpha\in(0,1)$. For $t=1-\alpha$, find $\sigma>0$ small enough and $k_{0}$ such that ($\star$) holds and such that the set $B(\sigma):=X\setminus C_{s}(\sigma)$ is not empty (the latter is possible by the condition on $X$). Then, for every $y\in X$ and every $k>k_{0}$,
(A.3) $\{j:0\leq j<k,f^{j}(y)\in B(\sigma)\}>\alpha k$
whenever $y,f(y),...,f^{k}(y)\notin C_{s}(\sigma)$. On the other hand, every $x\in B(\sigma)$ has a small neighborhood $U_{x}$ such that $f:U_{x}\to\mathbb{C}$ is a homeomorphism. Let $\delta>0$ be the Lebesgue number of the covering $(U_{x})$ of $B(\sigma)$. Then
(A.4) $x_{1},x_{2}\in B(\sigma),\ |x_{1}-x_{2}|<\delta\Rightarrow f(x_{1})\neq f(x_{2}).$
Now, take any $x\in B(\sigma)$. As $C_{s}(\sigma)$ is forward invariant, $B(\sigma)$ is backward invariant. It follows from the condition on $X$ that we can choose $x$ not to be in the forward orbit of any critical point of $f$.
We are in a position to repeat literally the end of the proof of [8] (where $B=B(\sigma)$).
A set $P\subset X$ is called $\delta$-separated if $|z_{1}-z_{2}|>\delta$ for all distinct $z_{1},z_{2}\in P$ and $P$ is $(k,\delta)$ separated if $\max\{|f^{j}(z_{1})-f^{j}(z_{2})|>\delta,j=0,...,k-1\}$ for all distinct points $z_{1},z_{2}\in P$ (in particular, any singleton is $(k,\delta)$-separated). Consider the backward orbit $BO(x)=\cup_{k=0}^{\infty}f^{-k}(x)$ of $x$. Then $f^{-1}(y)=N$ for all $y\in BO(x)$. For every $y\in BO(x)$ there exists a $\delta$-separated set $P(y)\subset f^{-1}(y)$ such that either $P(y)=\{z\}$ for some $z\notin B(\sigma)$ or $P(y)\subset B(\sigma)$ and #P(y)=N\#P(y)=N. Fix $k>k_{0}$ and define by induction:
$Q_{0}=\{x\};\ \ Q_{j+1}=\cup_{y\in Q_{j}}P(y),\ \ j=0,1,...,k-1.$
Then $Q_{k}$ is $(k,\delta)$-separated (induction in j=0,,kj=0,...,k). Let us show that $Q_{k}\geq N^{\alpha k}$. For $y\in Q_{j}$, let $Gen(y,j)=\#\{i:0\leq i<j,f^{i}(y)\in B(\sigma)\}$. Let $m=[\alpha k]+1$ (where $[T]$ is the entire part of $T>0$). Then, by (A.3), there are $N^{m}$ pairs $(y,j)$ such that $y\in B(\sigma)\cap Q_{j}$ and $Gen(y,j)=m$. For each such a pair there exists $z\in Q_{k}$ for which $f^{k-j}(z)=y$ and such points are distinct for distinct pairs. Therefore, $Q_{k}\geq N^{m}\geq N^{\alpha k}$. We have shown that for every $\alpha\in(0,1)$ there exist $k_{0}$ and $\delta>0$ such that the maximal cardinality $r_{k}(f,\delta)$ of a $(k,\delta)$-separated subset of $X$ is bigger than or equal to $N^{\alpha k}$, for all $k>k_{0}$. Thus $h_{top}(f)=\lim_{\delta\to 0}\limsup_{k\to\infty}\frac{1}{k}\log r_{k}(f,\delta)\geq\alpha\log N$, for all $\alpha\in(0,1)$. We are done.
(2) Suppose that $W$ is simply connected and $X\subset W^{\prime}\cap W$ consists of cycles of critical points of $f$. Then $f^{-1}(x)=1$ for all $x\in X$ which, along with the Riemann-Hurwitz formula for the map $f:U\to W$ where $U$ is a component of $W^{\prime}$ yield that $W^{\prime}$ consists of a single component and $X$ is just a point (which is then a fixed critical point of $f:W^{\prime}\to W$ of the order $d$).
Corollary A.1.
Under the conditions of Theorem A.1, the logarithmic capacity of the compact set $X$ is positive.
Proof.
By the Variational Principle, $\log N\leq h_{top}(f)=\sup h_{\mu}(f)$, where supremum is taken over the set $M(f)$ of all $f$-invariant Borel probability measures on $X$ and $h_{\mu}(f)$ is the entropy of $f$ w.r.t. $\mu$, see e.g. [14]. In particular, there exists an ergodic $\mu\in M(f)$ with $h_{\mu}(f)>0$. This implies that the Lyapunov exponent $\chi_{\mu}(f):=\int\log|f^{\prime}|d\mu>0$. Therefore, $HD(\mu)=h_{\mu}(f)/\chi_{\mu}(f)>0$, where $HD(\mu)$ is the infimum of Hausdorff dimensions over all Borel $Y\subset X$ such that $\mu(Y)=1$. Hence, Hausdorff dimension of $X$ is positive. In turn, this implies (see e.g. [17]) that the logarithmic capacity of $X$ is positive as well. ∎
Remark A.2.
By [14], Theorem 11.6.1, any such ergodic $\mu\in M(f)$ is approximated by measures supported on Cantor expanding repellers of $f$. In particular, there exist a lot of periodic orbits in any neighborhood of $X$. However, in general they are off $X$.
Appendix B Uniformization
Theorem B.1.
(see e.g. [17]) Let $X$ be a full compact set in $\mathbb{C}$ that contains at least $3$ points. Let $\pi:\mathbb{D}=\{|w|<1\}\to\hat{C}\setminus X$ be a universal holomorphic covering map and $\Gamma=\{\gamma\}$ be the corresponding Fuchsian group of Mobius transformations of $\mathbb{D}$ without elliptic elements such that $\mathbb{C}\setminus X\simeq\mathbb{D}/\Gamma$. Normalize $\pi$ such that $\pi(0)=\infty$. The following conditions are equivalent:
  1. (1)
    logarithmic capacity $c(X)$ of $X$ is positive,
  2. (2)
    $\Gamma$ is of convergence type, that is
    $\sum_{\gamma\in\Gamma}(1-|\gamma(w)|)<\infty$
    for some (hence, any) $w\in\mathbb{D}$. In this case, the infinite Blaschke product
    $B(w)=\Pi_{\gamma\in\Gamma}\frac{|\gamma(0)|}{\gamma(0)}\frac{w-\gamma(0)}{1-\overline{\gamma(0)}w}$
    is a holomorphic function in $\mathbb{D}$ such that $|B\circ\gamma(w)|=|B(w)|$ for all $w\in\mathbb{D}$, $\gamma\in\Gamma$, $|B(w)|<1$ for all $w\in\mathbb{D}$ and, for almost every $w_{0}\in\partial\mathbb{D}$, the angular limit $\lim|B(w)|=1$ as $w\to w_{0}$ exists.
    Moreover, projecting the function $\log|B|$ to $\mathbb{C}\setminus X$ we obtain Green’s function $G_{X}$ of $\mathbb{C}\setminus X$, that is,
    $G_{X}(z)=-\log|B(w)|,$
    for some (hence, every) $w\in\mathbb{D}\setminus\{\gamma(0):\gamma\in\Gamma\}$, such that $z=\pi(w)$. Then $G_{X}$ is well-defined everywhere positive harmonic function in $\mathbb{C}\setminus X$, $G(z)-\log|z|\to-\log(c(X))$ as $z\to\infty$ and, for almost every $w_{0}\in\partial\mathbb{D}$, the following angular limit exists:
    (B.1) $\lim_{w\to w_{0}}G_{X}(\pi(w))=0.$
Definition B.1.
For an angle $\beta\in(0,\pi/2)$, $r\in(0,\sin\beta)$, and a point $a\in\partial\mathbb{D}$, let $\Delta_{a}(\beta,r)$ be a connected component of the set $\{z:r<|z|<1,|\arg(z-a)|<\beta\}$ which contains $a$ in its boundary.
In other words, $\Delta_{a}(\beta,r)$ is an open "triangle" at a vertex $a$ bounded by an arc of the circle $|z|<r$ and segments of two rays through $a$ which form the angle $\beta$ with the radius $[0,a]$ of the unit circle.
Recall that $|F|$ denotes the Lebesgue 1-dim measure (length) of a measurable subset $F\subset\partial\mathbb{D}$
We have:
Corollary B.2.
Let $X$ be a compact with positive logarithmic capacity. Then for every $\varepsilon>0$ there exists a domain $W\subset\mathbb{D}$ as follows:
  1. (1)
    $W$ is simply connected,
  2. (2)
    $\partial W$ is a rectifiable Jordan curve,
  3. (3)
    there exists a point $w_{0}\in\partial W$ such that $|B(w)|>|B(w_{0})|$ for all $w\in W$ and $\pi(w)=\pi(w_{0})$ for finitely many $w\in\partial W$,
  4. (4)
    $|\partial W\cap\partial\mathbb{D}|>0$,
  5. (5)
    $\pi(W)$ lies in the $\varepsilon$-neighborhood of $X$,
  6. (6)
    $\pi(w)\to X$ uniformly as $w\to\partial\mathbb{D}$ inside of $W$.
Proof.
Let $X_{\varepsilon}=\{z\in\mathbb{C}\setminus X:d(z,X)<\varepsilon\}$. Define
$l_{\varepsilon}=\inf\{G_{X}(z):z\in\mathbb{C}\setminus X_{\varepsilon}\}.$
Then $l_{\varepsilon}>0$ and if $G_{X}(z)<l_{\varepsilon}$ for some $z\in\mathbb{C}\setminus X$, then $z\in X_{\varepsilon}$.
Let $b=\exp(-l_{\varepsilon})$. Then $|B(w)|>b$ implies $\pi(w)\in X_{\varepsilon}$.
Now we repeat partly a Lusin-Privalov construction (as in the proof of Privalov’s theorem, e.g. [17]). Fix an angle $\beta\in(0,\pi/2)$ and given $a\in\partial\mathbb{D}$ and $n\in\mathbb{N}$, such that $1-1/n<\sin\beta$ define
$H_{n}(a)=\inf\{|B(w)|:w\in\Delta_{a}(\beta,1-1/n)\}.$
By Theorem B.1,
$\lim_{n\to\infty}H_{n}(a)=1$
for a.e. $a\in\partial\mathbb{D}$. Hence, by Egorov’s theorem, there exists a closed subset $F\subset\partial\mathbb{D}$ of positive measure such that $H_{n}(a)\to 1$ as $n\to\infty$ uniformly in $a\in F$. In particular, for some $n_{0}\in\mathbb{N}$,
$|B(w)|>b$
whenever
$w\in W_{*}=\cup_{a\in F}\Delta_{a}(\beta,1-1/n_{0}).$
One can assume that $F$ lies in an arc of $\partial\mathbb{D}$ of small enough length and $n_{0}$ is big enough so that the part of the boundary of $W_{*}$ which lies on the circle $|w|=1-1/n_{0}$ is a proper part of this circle. Note that $\partial W_{*}\cap\partial\mathbb{D}=F$. Hence, there exists a connected component $W$ of the (open) set $W_{*}$ (note that $W=W_{*}$ if $W_{*}$ is connected itself) and a closed subset $F_{0}$ of $F$ of positive measure as follows:
$W=\cup_{a\in F_{0}}\Delta_{a}(\beta,1-1/n_{0}),$
$|B(w)|>b$ for every $w\in W$ and $|B(w)|\to 1$ as $w\to F_{0}$ uniformly in $W$. Then properties (4)-(6) hold for $W$. For the property (3), let $w_{0}\in\overline{W}$ be so that $|B(w_{0})|=\inf\{|B(w)|:w\in\overline{W}\setminus\partial\mathbb{D}\}$. Then $w_{0}\in\partial W$ and $|B(w)|>|B(w_{0})|$, by the Minimum principle and there could be at most finitely many $w_{n}\in\partial W$ with $\pi(w_{n})=\pi(w_{0})$ as otherwise $w_{n}\to\partial W\cap\partial\mathbb{D}$ which is impossible as then $|B(w_{0})|=|B(w_{n})|\to 1$. Let’s show that (1)-(2) hold too. By the construction, $\partial W\cap\{|w|=1-1/n_{0}\}$ is a proper arc $\{(1-1/n_{0})\exp{it}:t_{-}\leq t\leq t_{+}\}$ of the circle $|w|=1-1/n_{0}$. Then the closure of $W$ is a subset of $R:=\{w=r\exp(it):1-1/n_{0}\leq r\leq 1,t_{-}\leq t\leq t_{+}\}$. Therefore, there exists a neighborhood $R^{+}$ of $R$ and a diffeomorphic in $R^{+}$ branch $\bf Log$} of the map $w\mapsto i\overline{\log(w)}$ such that (slightly abusing notation and replacing $W$ by ${\bf Log}(W)$) one can assume that $F_{0}$ is a compact subset of $\mathbb{R}$ and
$W=\{z=x+iy:0<y<h,|\arg(z-x_{0})|<\pi/2-\beta\mbox{ for some }x_{0}\in F_{0}\},$
where $h=-\log(1-1/n_{0})$. Then it is easy to see that the boundary of $W$ is connected, which implies that $W$ is simply connected. Now, if $I=[x_{1},x_{2}]\subset\mathbb{R}$ is the smallest closed interval containing $F_{0}$ and $I\setminus F_{0}=\cup_{j=1}^{\infty}I_{j}$ where $I_{j}$ are pairwise disjoint open intervals, then the boundary of $W$ consists of: the "upper" boundary interval $\{x+ih:x_{1}-h\tan(\beta)\leq x\leq x_{2}+h\tan(\beta)\}$, "left" and "right" intervals $x_{1}-h\tan(\beta)+ih,x_{1}]$ and $x_{2},x_{2}+h\tan(\beta)+ih]$, the compact set $F_{0}\subset\mathbb{R}$ and a "saw"-like curve consisting of a sequence of two sides of equilateral triangles with bases $I_{j}$ whose total length is bounded by $\sum_{j}|I_{j}|/\sin(\beta)=(|I|-|F_{0}|)/\sin(\beta)$. Thus $\partial W$ is a rectifiable curve.
Appendix C Boundary distortion of sets under univalent self-maps of a disk
Let $f:\mathbb{D}\to\mathbb{D}$ be a univalent map of the unit disk into itself. The following are known theorems, see [11].
  • For a.e. $\zeta\in S^{1}=\partial\mathbb{D}$, $f(z)$ has an angular limit $f(\zeta)$.
  • $f$ has a finite angular derivative at $\zeta\in S^{1}$ if and only if the function $f^{\prime}(z)$ has an angular limit $f^{\prime}(\zeta)$ at $\zeta$.
  • If, for some $\zeta\in S^{1}$, $f(\zeta)\in S^{1}$, then (the angular derivative) $f^{\prime}(\zeta)$ exists and $1\leq\frac{\zeta}{f(\zeta)}f^{\prime}(\zeta)\leq\infty$ (the Wolff-Denjoy Theorem).
  • If $E\subset S^{1}$ is measurable and $f(E)\subset S^{1}$, then $f^{\prime}(\zeta)$ is finite for a.e. $\zeta\in E$ (a corollary from the McMillan Twist Theorem).
In the next lemma, the inequality (C.1) of part 1 is proved in [10] [^1] Although not explicitly stated in [10], the inequality (C.1) is attained as the limit case, as $\to 1, zz\in\mathbb{R}$, in the inequality at the top of p.259 assuming w.l.o.g. that $\zeta=1$., [15], [12]. Part 2 is an easy consequence, see below.
Lemma C.1.
Let $f:\mathbb{D}\to\mathbb{D}$ be univalent. For some $\zeta\in S^{1}$, suppose that the angular limit $f(\zeta)$ exists and $f(\zeta)\in S^{1}$. Assume that the angular derivative $f^{\prime}(\zeta)$ is finite.
  1. (b1)
    If $f(0)=0$, then
    (C.1) $|f^{\prime}(0)|^{-1/2}\leq|f^{\prime}(\zeta)|.$
  2. (b2)
    in general,
    (C.2) $\frac{(1-|f(0)|)^{3/2}}{(1+|f(0)|)^{1/2}}|f^{\prime}(0)|^{-1/2}\leq|f^{\prime}(\zeta)|.$
Proof of (b2): let $g(z)=\frac{f(z)-A}{1-\bar{A}f(z)}$ where $A=f(0)$. As $g^{\prime}(z)=\frac{f^{\prime}(z)(1-|A|^{2})}{(1-\bar{A}f(z))^{2}}$, then $g^{\prime}(0)=\frac{f^{\prime}(0)}{1-|A|^{2}}$ and $|g^{\prime}(\zeta)|=\frac{|f^{\prime}(\zeta)|(1-|A|^{2})}{|1-\bar{A}f(\zeta)|^{2}}\leq\frac{|f^{\prime}(\zeta)|(1-|A|^{2})}{(1-|A|)^{2}}=\frac{|f^{\prime}(\zeta)|(1+|A|)}{(1-|A|)}$. By (C.1), this implies (C.2).
The following lemma and its corollary state that, if $E,f(E)\subset S^{1}$, then the size of $E$ is small provided $|f^{\prime}(0)|$ is small and $f(0)$ is away from the unit circle. Exact bounds are obtained in [15], under the assumption that $E$ is contained in an open set $E_{0}\subset S^{1}$ such that $f(E_{0})\subset S^{1}$ (and $f(0)=0$).
Lemma C.2.
Let $f:\mathbb{D}\to\mathbb{D}$ be univalent. Let $E\subset S^{1}$ be a measurable set such that $f(E)\subset S^{1}$, and $|E|>0$. Then for every $\epsilon>0$ there exists a closed set $L\subset E$ such that $|L|>|E|-\epsilon$, $f(L)$ is measurable and
(C.3) $|f(L)|\geq|L|\sqrt{\frac{(1-|f(0)|)^{3}}{(1+|f(0)|)|f^{\prime}(0)|}}.$
Proof.
As $f^{\prime}(\zeta)$ exists and is finite for a.e. $\zeta\in E$, and $f^{\prime}(\zeta)=\lim_{n\to\infty}f^{\prime}((1-1/n)\zeta)$ where $f^{\prime}$ is continuous in $\mathbb{D}$, then $f^{\prime}:E\to\mathbb{C}$ is measurable. By Lusin’s theorem, there exists a closed subset $E^{\prime}$ of $E$ such that $|E^{\prime}|>|E|-\epsilon/2$ and $f^{\prime}:E^{\prime}\to\mathbb{C}$ is (uniformly) continuous. Fix an angle $\beta\in(0,\pi/2)$ and a radius $r\in(0,\sin\beta)$.
Given $X\subset S^{1}$, let
$W_{X}=\cup_{a\in X}\Delta_{a}(\beta,r)$
(see Definition B.1)). For every $n$ big enough, define $M_{n}:E^{\prime}\to\mathbb{R}$ as follows:
$M_{n}(a)=\sup\{|f^{\prime}(z)-f^{\prime}(a)|:z\in\Delta_{a}(\beta,1-1/n).$
Since $f^{\prime}(z)\to f^{\prime}(a)$ as $z\to a$ while $|\arg(z-a)|<\beta$, $M_{n}(a)\to 0$ for every $a\in E^{\prime}$. By Egorov’s theorem, there exists a closed set $L\subset L^{\prime}$, $|L|>|L^{\prime}|-\epsilon/2$, such that $M_{n}\to 0$ uniformly on $L$. (Note that $|E|>|E|-\epsilon$.) Consider $f^{\prime}$ on $W_{L}$. We show that $f^{\prime}(z)\to f^{\prime}(a)$ as $z\to a$ in $W_{L}$ uniformly in $a\in L$. Indeed, given a small $t>0$, choose a small $\delta_{0}>0$ such that $|f^{\prime}(b)-f^{\prime}(a)|<t$ whenever $a,b\in L$ and $|b-a|<\delta_{0}$, and choose $N$ such that $M_{N}(b)<t/2$ for all $b\in L$. Now, let $\delta=\min\{1/N,\delta_{0}/(2\cos\beta)\}$. It is chosen so that, for any $z\in B(1,\delta)\cap\mathbb{D}$, if $z\in\Delta_{c}(\beta,r)$ for some $c\in S^{1}$ then $|c-1|<\delta_{0}$.
If now $|z-a|<\delta$ for some $a\in L$, $z\in W_{L}$, then, by the definition of $W_{L}$, $z\in\Delta_{b}(\beta,r)$ for some $b\in L$, on the other hand, $|b-a|<\delta_{0}$, by the choice of $\delta$. Therefore,
$|f^{\prime}(z)-f^{\prime}(a)|\leq|f^{\prime}(z)-f^{\prime}(b)|+|f^{\prime}(b)-f^{\prime}(a)|<M_{N}(b)+t/2<t/2+t/2=t.$
Thus $f^{\prime}(z)\to f^{\prime}(a)$ as $z\to a$ for $z\in W_{L}$, uniformly in $a\in L$. Finally, as $f^{\prime}$ is continuous in $\mathbb{D}$, we conclude that $f^{\prime}$ is a continuous (in particular, bounded) function in the closure $\overline{W_{L}}$ of $W_{L}$.
Since $L=\overline{W_{L}}\cap S^{1}$, it is enough to prove (C.3) with $L$ replaced by $F\subset L$ such that $W_{F}$ is a component of the (open) set $W_{E}$. Fix such an $F$. $W_{F}$ is a simply connected domain, hence, there exists a conformal isomorphism $R:\mathbb{D}\to W_{F}$. As $W_{F}$ is bounded by a rectifiable Jordan curve, by Riesz-Privalov’s theorem, the derivative $R^{\prime}$ belongs in the Hardy space $H^{1}(\mathbb{D})$. Hence, $R^{-1}(F)$ is measurable. Consider a univalent map $g:=f\circ R:\mathbb{D}\to\mathbb{D}$. As $f^{\prime}$ is bounded in $W_{F}$, $g^{\prime}\in H^{1}(\mathbb{D})$ as well and the set $g(R^{-1}(F))$ is measurable. By Riesz-Privalov’s theorem, we can write:
$|f(F)|=|g(R^{-1}(F))|=\int_{R^{-1}(F)}|g^{\prime}(w)||dw|=\int_{R^{-1}(F)}|f^{\prime}(R(w))||R^{\prime}(w)||dw|=$
$\int_{F}|f^{\prime}(z)||dz|\geq|F|\sqrt{\frac{(1-|f(0)|)^{3}}{(1+|f(0)|)|f^{\prime}(0)|}},$
where the latter inequality follows from Lemma C.1.
Corollary C.3.
Let $f:\mathbb{D}\to\mathbb{D}$ be univalent. Let $E\subset S^{1}$ be a measurable set such that $f(E)\subset S^{1}$, and $|E|>0$. Then
(C.4) $|E|\sqrt{\frac{(1-|f(0)|)^{3}}{(1+|f(0)|)|f^{\prime}(0)|}}\leq 2\pi.$